### v1

# Efficient Bayesian Additive Regression Models For Microbiome Studies

Abstract
Statistical analysis of microbiome data is challenging. Bayesian multinomial logistic-normal (MLN) models have gained popularity due to their ability to account for the count compositional nature of these data. However, these models are often computationally intractable to infer. Recently, we developed a computationally efficient and accurate approach to inferring MLN models with a Marginally Latent Matrix-T Process (MLTP) form: MLN-MLTPs. Our approach is based on a novel sampler with a marginal Laplace approximation – called the Collapse-Uncollapse (CU) sampler. However, existing work with MLTPs has been limited to linear models or models of a single non-linear process. Moreover, existing methods lack an efficient means of estimating model hyperparameters. This article addresses both deficiencies. We introduce a new class of MLN Additive Gaussian Process models (MultiAddGPs) for deconvolution of overlapping linear and non-linear processes. We show that MultiAddGPs are examples of MLN-MLTPs and derive an efficient CU sampler for this model class. Moreover, we derive efficient Maximum Marginal Likelihood estimation for hyperparameters in MLTP models by taking advantage of Laplace approximation in the CU sampler. We demonstrate our approach using simulated and real data studies. Our models produce novel biological insights from a previously published artificial gut study.
Additive Regression,
Multinomial Logistic-Normal Models,
Bayesian Inference,
Microbiome Data,
keywords:
[…]
, and
1 Introduction
Dysregulation of human-, animal-, and even plant-associated microbial communities (microbiota) are known to cause disease (Ballen et al., 2016; Frati et al., 2018; Holleran et al., 2018; Sharon et al., 2019; Gao et al., 2021). In humans, alterations of microbiota play a causal role in obesity (Tilg et al., 2011; Ley, 2010), inflammatory bowel disease (Honda and Littman, 2012; Glassner, Abraham and Quigley, 2020; Kostic, Xavier and Gevers, 2014), and even cancer (Schwabe and Jobin, 2013; Helmink et al., 2019). As a result, many researchers study how dietary, host physiologic, and environmental factors influence the relative abundance of different bacterial taxa in microbiota. These factors can have linear or non-linear effects on community structure (Cheng et al., 2019; Sankaran and Jeganathan, 2024; Schwager et al., 2017). Even within closed in vitro systems, microbiota can demonstrate non-linear temporal variation (Silverman et al., 2018). Overall, flexible statistical methods are needed to disentangle linear and non-linear effects on microbiota.
Beyond the biological complexity of microbiota, limitations of the measurement process further complicate analyses. These data are typically represented as a $D\times N$ count table $Y$ with elements $Y_{dn}$ denoting the number of DNA molecules from taxon $d$ observed (sequenced) in sample $n$. Critically, the counts $Y_{dn}$ do not represent the true abundance of taxon $d$ in the biological system (e.g., colon) from which the sample $n$ was obtained. Instead, the counts are the results of a random sample of that pool of microbes. Due to limitations of the measurement process, the size of that sample (the sequencing depth; $\sum_{d=1}^{D}Y_{dn}$) is typically arbitrary and unrelated to the total microbial load in the system (Vandeputte et al., 2017). As a result, many authors call these data compositional, reflecting the idea that the data only provide information about the relative abundances of the different taxa within each community (Gloor et al., 2017; McGregor et al., 2023; Mao and Ma, 2022). However unlike traditional compositional data (e.g., continuous simplex valued data), these data are zero-laden counts. The presence of counting noise and zeros limits standard approaches to compositional data analysis, which typically involve analyses of log-ratio transformed data (Kuczynski et al., 2012; Silverman et al., 2018, 2017, 2020; Kaul et al., 2017; Li, 2015; Gloor et al., 2016).
Many researchers have turned to Bayesian Multinomial Logistic Normal (MLN) models to address the measurement process’s challenges (Silverman et al., 2021; Äijö, Müller and Bonneau, 2018; Grantham et al., 2020; Silverman et al., 2018, 2022). The multinomial is used to model uncertainty due to random counting, while the logistic normal captures the extra-multinomial variability typically seen in these data (Silverman et al., 2022). More conceptually, Bayesian MLN models allow researchers to model community composition as a latent simplicial vector that is informed by the observed count data. Moreover, unlike the more well-known Dirichlet distribution, the logistic-normal has a rich covariance structure which allows modeling both positive and negative covariation between taxa (Aitchison and Shen, 1980; Silverman et al., 2018). The logistic-normal is also self-conjugate (as it is multivariate normal under a suitable log-ratio transformation), allowing for a wide variety of models to be built in the latent simplex space. However, the multinomial and the logistic-normal are not conjugate, making inference of these models computationally challenging or even intractable.
Recent advances have made Bayesian MLN models practical for microbiota analyses. Early Bayesian MLN models were inferred via Metropolis-within-Gibbs and only scaled to a handful of dimensions (Cargnoni, Müller and West, 1997). Pólya-Gamma data augmentation has been used successfully in some instances (Glynn et al., 2019); however, it is still too computationally intensive for widespread use due to the inability to perform blocked Gibbs updates of Pólya-Gamma random variables while maintaining the logistic-normal form (Linderman, Johnson and Adams, 2015). Subsequently, researchers successfully used Hamiltonian Monte Carlo (Äijö, Müller and Bonneau, 2018; Grantham et al., 2020; Silverman et al., 2018). Yet scalability still limits those methods. For example, the sampler used in Silverman et al. (2018) took more than four hours to analyze a dataset of approximately one thousand samples, yet only ten taxa. More recently, we proved that a wide variety of Bayesian MLN models, including generalized linear models and generalized Gaussian process regression models, share a common marginal form called a Latent Matrix-T Process (LTP) (Silverman et al., 2022). We showed that a Laplace approximation to this marginal form was extremely accurate, leading to an efficient and accurate approximate inference procedure called the Collapse-Uncollapse (CU) sampler. Our result demonstrated that this approach is often 4-5 orders of magnitude faster than HMC-based methods with minimal error in posterior calculations (Silverman et al., 2022).
Despite these advances, there remains a dearth of tools for disentangling the effects of multiple measured factors on microbiota. Recently, Cheng et al. (2019) proposed an additive Gaussian process framework to address this need. Yet their approach assumed the data was transformed Gaussian, ignoring count compositional nature of these data. Moreover, our prior work with MLTPs was limited to factors that have a linear effect on microbial composition (generalized linear models) or a single factor that had a nonlinear effect (generalized Gaussian process regression models). Finally, our prior work did not address the problem of hyperparameter selection, which is often critical in both linear and nonlinear modeling (e.g., selection of kernel parameters).
This article addresses the limitations of prior methods and develops a flexible, and computationally efficient approach to disentangling both linear and nonlinear effects on microbiota. As in Cheng et al. (2019), our approach is based on a class of additive Gaussian process regression models. Unlike Cheng et al. (2019), we do not assume that the data is transformed Gaussian and instead prove that Bayesian MLN Additive Gaussian Process Models are also MLTPs. Using those results, we extend the CU sampler to this class of models. Moreover, we extend the results of Silverman et al. (2022) by developing an efficient maximum marginal likelihood estimator for model hyperparameters. Through simulation studies, we validate our methods; we find posterior intervals from our models often cover the truth ten to a hundred times more often than those of transformed Gaussian models. Finally, we apply our models to a previously published longitudinal in vitro gut microbiome study which studied the effects of starvation on microbial communities (Silverman et al., 2018). Our models are able to provide the first estimates of the impact of starvation over time within that study.
We organize this article as follows. Section 2 provides a review of the Marginally LTP class of Bayesian MLN models, as well as a review of the Collapse-Uncollapse sampler with marginal Laplace approximation. Section 3 presents a multinomial logistic-normal generalized additive Gaussian process regression (MultiAddGP) model and proves that this model is part of the Marginally LTP class. We also derive a computationally efficient approach to hyperparameter estimation in MLTP models via maximum marginal likelihood. Sections 4 and 5 demonstrate our approach through application to both simulated and real microbiome data. Finally, we conclude with a discussion in Section 6.
Notation
In this article, we denote matrix and vector dimensions with unbolded uppercase letters (e.g., $N$), matrices and matrix-valued functions using bold uppercase symbols (e.g., $\boldsymbol{X}$), vectors and vector-valued functions with bold lowercase symbols (e.g., $\mathbf{x}$), and scalars and scalar functions as unbolded lowercase symbols (e.g., $x$). For matrices, we index specific rows as $\mathbf{X}_{d\cdot}$ and columns as $\mathbf{X}_{\cdot n}$. We denote vector-valued stochastic processes using the same notation as matrices (e.g. $\boldsymbol{Y}$) since, in practice, we only evaluate these at a finite number of test points.
2 Review of Marginally Latent Matrix-T Processes and the CU Sampler
We describe the class of Marginally Latent Matrix-T Processes (MLTPs) by sequentially generalizing from Matrix-T Processes to Latent Matrix-T Processes (LTPs) and finally MLTPs. We then describe the subset of Bayesian Multinomial Logistic-Normal MLTPs (MLN-MLTPs) before reviewing the inference of this class of models.
2.1 Defining Marginally Latent Matrix-T Processes (MLTPs)
Just as Gaussian processes can be defined based on the marginal properties of the multivariate normal, matrix normal processes and Matrix-T processes can be defined by the marginal properties of the matrix normal and matrix-T distributions (Silverman et al., 2022). Matrix-T processes generalize Student-T processes and Gaussian processes (Silverman et al., 2022).
Definition 2.1 (Matrix-T Process).
A stochastic process $\mathbf{Y}\sim TP(\nu,\mathbf{M},\mathbf{V},\mathbf{A})$ defined on the set $\mathcal{W}=\mathcal{W}^{(1)}\times\mathcal{W}^{(2)}$ is a matrix-T process if $\mathbf{Y}$ evaluated on any two finite subsets $\mathcal{X}^{(1)}\subset\mathcal{W}^{(1)}$ and $\mathcal{X}^{(2)}\subset\mathcal{W}^{(2)}$ is a random matrix $\mathbf{Y}$ of dimension $|\mathcal{X}^{(1)}|\times|\mathcal{X}^{(2)}|$ that follows a matrix-T distribution: $\mathbf{Y}\sim T(\nu,\mathbf{M,V,A})$. $\nu$ is a scalar value strictly greater than zero. Let $x_{i}^{(1)},x_{j}^{(1)}\in\mathcal{X}^{(1)}$ and $x_{i}^{(2)},x_{j}^{(2)}\in\mathcal{X}^{(2)}$. $\mathbf{M}_{ij}=\mathbf{M}(x_{i}^{(1)},x_{j}^{(2)})$ is the matrix function representing the mean, and $\mathbf{V}_{ij}=\mathbf{V}(x_{i}^{(1)},x_{j}^{(1)})$ and $\mathbf{A}_{ij}=\mathbf{A}(x_{i}^{(2)},x_{j}^{(2)})$ are kernel functions.
Latent Matrix-T Processes (LTP) generalize Matrix-T processes. $\mathbf{Y}$ is said to be an LTP if
$\mathbf{Y}$ $\sim g(\mathbf{\Pi},\lambda)$
$\mathbf{\Pi}$ $=\phi^{-1}(\mathbf{H})$
$\mathbf{H}$ $\sim TP(\nu,\mathbf{M},\mathbf{V},\mathbf{A}).$
where $g$ is any distribution depending on parameters $\mathbf{\Pi}$ as well as hyperparameters $\lambda$ and $\phi$ is a known transform. LTPs can alternatively be written as a joint model $p(\mathbf{Y},\mathbf{H})$.
A stochastic process $\mathbf{Y}$ is Marginally LTP (MLTP) if it can be described by a joint distribution $p(\mathbf{Y},\mathbf{H},\mathbf{\Phi})$ with a marginal $p(\mathbf{Y},\mathbf{H})$ that is a LTP. Silverman et al. (2022) showed that a wide variety of linear, dynamic linear, and non-linear regression models are MLTP. In Section 3, we show that our proposed class of generalized additive Gaussian process regression models are MLTP as well.
2.2 Bayesian Multinomial Logistic Normal MLTPs (MLN-MLTPs)
Bayesian Multinomial Logistic Normal MLTPs (MLN-MLTPs) are a subtype of MLTPs that are particularly useful for the analysis of microbiome data. In MLN-MLTPs, the distribution $g$ is a product multinomial: $p(\mathbf{Y}_{\cdot 1},\dots,\mathbf{Y}_{\cdot N})\sim\prod_{n=1}^{N}\text{Multinomial}(\boldsymbol{\Pi}_{\cdot n})$ and the transform $\phi$ is an invertible log-ratio transform from the $D$-dimensional simplex to $D-1$ dimentional real-space: $\mathbf{H}_{\cdot n}=\phi(\boldsymbol{\Pi}_{\cdot n}\in\mathbb{S}^{D})\in\mathbb{R}^{D-1}$. Canonically, we used the following Additive Log-Ratio (ALR) transform which takes the $D$-th taxa as a reference:
$\mathbf{H}_{.n}=\phi(\boldsymbol{\Pi}_{.n})=\left\{\log\left(\frac{\pi_{1n}}{\pi_{Dn}}\right),\ldots,\log\left(\frac{\pi_{(D-1)n}}{\pi_{Dn}}\right)\right\}^{T}.$ (1)
We choose this transform for computational efficiency as discussed in Silverman et al. (2022). There is no loss in generality as posterior samples taken with respect to the $\text{ALR}_{D}$ coordinate system can be transformed into any other log-ratio coordinate system (Pawlowsky-Glahn, Egozcue and Tolosana-Delgado, 2015, Appendix A.3). For context, the $\text{ALR}_{D}$ transform is the inverse of the softmax transform.
2.3 Collapsed-Uncollapsed (CU) Sampler
The definition of MLTPs is key to efficient inference. If a model $p(\mathbf{Y},\mathbf{H},\mathbf{\Phi})$ has a closed-form marginal $p(\mathbf{Y},\mathbf{H})$ that is an LTP, then its closed form conditional $p(\mathbf{\Phi}\mid\mathbf{Y},\mathbf{H})$ likely exists. We call the marginal $p(\mathbf{Y},\mathbf{H})$ the collapsed form and $p(\mathbf{\Phi}\mid\mathbf{Y},\mathbf{H})$ the uncollapsed form. The posterior of an MLTP factors as
$p(\mathbf{H},\Phi\mid\mathbf{Y})=p(\mathbf{\Phi}\mid\mathbf{H},\mathbf{Y})p(\mathbf{H}\mid\mathbf{Y})$
with the uncollapsed form as the first term and the posterior of the collapsed form as the second. As the collapsed form is rarely conjugate, techniques such as MCMC can be used to obtain samples from it’s posterior. Then, conditioned on those samples, the uncollapsed form can be used to obtain samples from the joint posterior. Especially when $\boldsymbol{\Phi}$ is high-dimensional, this Collapse-Uncollapse sampler can be much more efficient than common alternatives (Silverman et al., 2022). Still, the most substantial enhancements occur when approximations to the collapsed form are considered.
We have developed a Laplace approximation for the collapsed form of MLTPs:
$p(vec(\mathbf{H})\mid\mathbf{Y})\approx N(vec(\hat{\mathbf{H}}),\nabla^{-2}[vec(\hat{\mathbf{H}})])$
where $\hat{\mathbf{H}}$ denotes the Maximum A Posteriori (MAP) estimate of the collapsed form and $\nabla^{-2}[vec(\hat{\mathbf{H}})]$ denotes the inverse Hessian of the collapsed form evaluated at the MAP estimate. To facilitate this approximation, we derived analytical results for the gradient and Hessian of the collapsed form (Silverman et al., 2022). Focusing on applications to MLN-MLTPs, we proved error bounds on the Laplace approximation and provided simulation and real analyses showing that the approximation was extremely accurate in the context of microbiome data analysis. Beyond the accuracy of posterior calculations, we showed that this CU sampler with the Laplace approximation (simply referred to as the CU sampler in the following text) was often 4-5 orders of magnitude faster than MCMC and 1-2 orders of magnitude faster than black-box variational inference while also being more accurate than the later. The CU sampler for MLN-MLTPs, along with uncollapse samplers for linear and non-linear regression models, is publicly available on CRAN as part of the fido software package.
3 Methods
3.1 Multinomial Logistic Normal Additive Gaussian Process Models (MultiAddGP)
To facilitate additive linear and non-linear modeling using Bayesian MLNs, this article introduces the class of Multinomial Logistic Normal Additive Gaussian Process Models (MultiAddGPs). Let $\mathbf{Y}_{\cdot n}$ denote a $D$-vector of observed data, $\mathbf{X}_{\cdot n}$ a $Q_{0}$-vector of covariates to model linearly, and each $\mathbf{Z}^{(k\in\{1,\dots,K\})}_{\cdot n}$ a $Q_{k}$-vector of covariates to be modeled with distinct non-linear functions. MultiAddGP models have the following form:
$\mathbf{Y}_{\cdot n}$ $\sim\text{Multinomial}(\boldsymbol{\Pi}_{\cdot n})$ (2)
$\boldsymbol{\Pi}_{\cdot n}$ $=\phi^{-1}(\mathbf{H}_{\cdot n})$ (3)
$\mathbf{H}_{\cdot n}$ $\sim N(\mathbf{F}_{\cdot n},\boldsymbol{\Sigma})$ (4)
$\mathbf{F}$ $=\mathbf{B}\mathbf{X}+\sum_{k=1}^{K}\mathbf{f}^{(k)}(\mathbf{Z}^{(k)})$ (5)
with priors $\mathbf{B}\sim N(\boldsymbol{\Theta}^{(0)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(0)})$, $\mathbf{f}^{(k)}\sim\text{GP}(\boldsymbol{\Theta}^{(k)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(k)})$, and $\boldsymbol{\Sigma}\sim\text{InvWishart}(\boldsymbol{\Xi},\zeta)$. As in prior sections, $\phi$ denotes any log-ratio transform from $\mathbb{S}^{D}$ to $\mathbb{R}^{D-1}$. $\boldsymbol{\Sigma}$ is a $(D-1)\times(D-1)$ covariance matrix. For the matrix-normal prior on the linear term, $\boldsymbol{\Theta}^{(0)}$ is the mean matrix and $\boldsymbol{\Gamma}^{(0)}$ is a $Q_{0}\times Q_{0}$ covariance matrix representing covariance in the parameters of the $Q_{0}$ covariates. The terms $\boldsymbol{\Theta}^{(k)}$ and $\boldsymbol{\Gamma}^{(k)}$ in the $K$ matrix-normal process priors echo their linear counterparts but are functions (e.g., mean and kernel functions) rather than fixed dimensional matrices. Additionally, we permit hyperparameters $\Omega$ in any part of the model including but not limited to hyperparameters in the kernel functions (e.g., bandwidth parameters) or hyperparameters in the mean functions. As we will show through simulated and real data analyses in Sections 4 and 5, this is a very flexible form of model which can be used in a wide range of additive linear and non-linear modeling tasks.
3.2 Posterior Estimation in MultiAddGPs
We use MLTP theory to sample from the posterior of MultiAddGPs: $p(\mathbf{H},\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)},\boldsymbol{\Sigma}\mid\mathbf{Y})$. To connect to that theory we first prove that MultiAddGPs are MLN-MLTP’s with $\boldsymbol{\Phi}=\{\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)},\boldsymbol{\Sigma}\}$. We derive the collapsed form and then derive an efficient algorithm for sampling from the uncollapsed form. First however, we must clarify the definition of $\mathbf{F},\mathbf{f}_{1},\dots,\mathbf{f}_{K}$.
Up to this point, we have not distinguished between the set of points $n\in\{1,\dots,N\}$ at which we have observed data $\mathbf{Y}$ and the potentially different set $n^{*}\in\{1,\dots,N^{*}\}$ at which we want to evaluate the functions $\mathbf{F},\mathbf{f}_{1},\dots,\mathbf{f}_{K}$. In what follows, we use the symbols $\mathbf{F},\mathbf{f}_{1},\dots,\text{ and }\mathbf{f}_{K}$ to denote the evaluation of corresponding infinite-dimensional functions at the set of evaluation points $\{1,\dots,N^{*}\}$, i.e., they are each $D\times N^{*}$-dimensional random matrices. In contrast, all other random matrices (e.g., $\mathbf{H}$) represent their corresponding infinite-dimensional analogues evaluated at the set of observed points ($\{1,\dots,N\}$).
In Supplementary Section 1 we prove that MultiAddGPs are MLN-MLTPs with the following parameters defining the collapsed form:
$\mathbf{M}$ $=\boldsymbol{\Theta}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Theta}^{(k)}(\mathbf{Z}^{(k)})$
$\mathbf{V}$ $=\boldsymbol{\Xi}$
$\mathbf{A}$ $=\mathbf{X}^{T}\boldsymbol{\Gamma}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}(\mathbf{Z}^{(k)})+\mathbf{I}$
$\mathbf{\nu}$ $=\mathbf{\zeta}$
[…]
Sampling from the uncollapsed form starts by obtaining samples from $p(\mathbf{F},\boldsymbol{\Sigma}\mid\mathbf{H})$. Conditioning on $\mathbf{H}$ and marginalizing over $\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}$ in the MultiAddGP model results in a Bayesian matrix-normal process model with likelihood $\mathbf{H}_{\cdot n}\sim N(\mathbf{F}_{\cdot n},\boldsymbol{\Sigma})$ and priors:
$\mathbf{F}$ $\sim\text{GP}\left(\boldsymbol{\Theta}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Theta}^{(k)}(\mathbf{Z}^{(k)}),\;\boldsymbol{\Sigma},\;\mathbf{X}^{T}\boldsymbol{\Gamma}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}(\mathbf{Z}^{(k)})\right)$
$\boldsymbol{\Sigma}$ $\sim\text{InvWishart}(\boldsymbol{\Xi},\zeta).$
This is the same model discussed in Silverman et al. (2022): samples from $p(\mathbf{F},\boldsymbol{\Sigma}\mid\mathbf{H})$ can be obtained via methods described in Appendix C of that article.
Finally, conditioned on samples of $\mathbf{F}$ and $\boldsymbol{\Sigma}$, we obtain samples of each $\mathbf{f}^{(k)}$. Inspired by the backfitting algorithm used for estimation in generalized additive models (Hastie and Tibshirani, 1990), we developed a backsampling algorithm for this task. The backsampling proceeds by iteratively sampling $p(\mathbf{B}\mid\mathbf{F})$, $p(\mathbf{f}_{1}\mid\mathbf{F},\mathbf{B})$, $p(\mathbf{f}_{2}\mid\mathbf{F},\mathbf{B},\mathbf{f}_{1})$, $\dots$, and then $p(\mathbf{f}^{(K)}\mid\mathbf{F},\mathbf{B},\mathbf{f}_{1},\dots,\mathbf{f}^{(K-1)})$. For brevity, we leave a description of this algorithm to Supplementary Sections 2 and 3.
3.3 Model Identification
[…]
$\tilde{\mathbf{f}}^{(k)}=\mathbf{f}^{(k)}(z^{(k)})-\frac{1}{n}\sum_{n=1}^{N}\mathbf{f}^{(k)}(z_{n}^{(k)}).$
3.4 Inferring Model Hyperparameters
Hyperparameter selection remains an outstanding problem in MLTPs (Silverman et al., 2022). Here we demonstrate how the same Laplace approximation used in the Collapse step of the CU sampler can be used to approximate marginal likelihoods and thereby estimate hyperparameters via Maximum Marginal Likelihood (MML) or penalized MML estimation.
Let $\Omega$ denote a set of hyperparameters in the joint MLTP distribution: $p(\mathbf{H},\Phi,\mathbf{Y}\mid\Omega)$. These can be parameters in a kernel, a mean function, or even in the log-ratio transformation $\phi$. The MML estimate for $\Omega$ is
[…]
where $\hat{\mathbf{H}}_{\Omega}$ denotes the MAP estimate for $\mathbf{H}$ evaluated in the observed set $\{1,\dots,N\}$ which depends on the hyperparameters $\Omega$. From a computational standpoint, this Laplace approximation to the marginal likelihood is essentially free given that we are already computing the MAP estimate and a factorization of the Hessian matrix as part of the CU sampler (Silverman et al., 2022).
While the MML estimate already provides a degree of penalization (Wood, 2011), it can be helpful to add an additional penalty term to the marginal likelihood when defining weakly identified MultiAddGPs. As we demonstrate in Section 5 and Supplementary Section 5, we often choose this penalty term as the log density of a prior for the hyperparameters.
In practice, we find that the CU sampler is efficient enough to make optimizing Equation (7) (with or without added penalization) practical. While we have found that various optimization routines can be used to perform the optimization, the problem is often non-convex, especially when estimating bandwidth parameters in kernel functions. For this reason, we generally recommend Bayesian optimization procedures which can optimize non-convex functions while simultaneously reducing the number of function evaluations compared to more typical optimizers (e.g., the L-BFGS optimizer). We have tested and found Bayesian optimization procedures generally work well when $\Omega$ is of low to moderate dimension (e.g., $\leq 10$). When $\Omega$ is higher dimensional, alternative procedures may be needed. In this study, Bayesian Optimization was performed using the ParBayesianOptimization package with the acquisition function set to the upper confidence bound, starting with 10 initial points and proceeding through 20 iterations with all other parameters set to their default values.
3.5 Software Availability
We extend the R package fido (Silverman et al. (2022)) to efficiently infer MultiAddGPs models. The fido package implements the extended CU sampler described above using optimized C++ code. MAP estimation of $\hat{H}$ is performed using the L-BFGS optimizer. Hyperparameter selection by MML is performed by maximizing the marginal likelihood using Bayesian optimization implemented in the ParBayesianOptimization package. All code required to reproduce the results of the this article is available at https://github.com/Silverman-Lab/MultiAddGPs.
4 Simulations
We simulated a suite of longitudinal studies of microbiota with varying numbers of taxa $D\in\{3,\dots,100\}$ and samples $N\in\{20,\dots,1000\}$. We simulated microbial composition influenced by batch effects, daily periodicity (e.g., circadian rhythm; Heddes et al. (2022)), and longer-term trends. Full simulation details are provided in Supplementary Section 4.
In Figure 1 we show a small simulation $D=4$ and $N=600$ for ease of visualization. We use $t_{n}$ to denote the time at which sample $n$ was obtained. For inference, we specify a MultiAddGP model $\mathbf{F}_{n}=\mathbf{b}_{0}+\mathbf{b}_{1}x^{(\text{batch})}_{n}+\mathbf{f}^{(\text{periodic})}(t_{n})+\mathbf{f}^{(\text{trend})}(t_{n})$ as follows. For covariates, we set $\mathbf{X}_{\cdot n}=[1\;x^{(\text{batch})}_{n}]^{T}$, $\mathbf{Z}^{(\text{periodic})}_{\cdot n}=t_{n}$, and $\mathbf{Z}^{(\text{trend})}_{\cdot n}=t_{n}$. For priors we set $\mathbf{B}=[\mathbf{b}_{0}=2.7;\mathbf{b_{1}}=1]$. Both $\mathbf{f}^{(\text{periodic})}$ and $\mathbf{f}^{(\text{trend})}$ were given matrix-normal process priors with mean function $\boldsymbol{\Theta}^{(k)}=\mathbf{0}$. A periodic kernel $K_{\text{period}}(t,t^{\prime})=\sigma_{\text{period}}\exp\left(-\frac{2\sin^{2}\left(\frac{\pi|t-t^{\prime}|}{p}\right)}{\rho_{\text{period}}^{2}}\right)$ was used in the prior for $\mathbf{f}^{(\text{periodic})}$ and a squared exponential $K_{\text{trend}}(t,t^{\prime})=\sigma_{\text{trend}}^{2}\exp(-\frac{(t-t^{\prime})^{2}}{2\rho_{\text{trend}}^{2}})$ was used for $\mathbf{f}^{(\text{trend})}$. Hyperparameters $\Omega=\{\sigma_{\text{period}},\rho_{\text{period}},p,\sigma_{\text{trend}},\rho_{\text{trend}}\}$ were selected using MML estimation.
For comparison, we created a nearly identical model that ignored uncertainty due to counting and assumed the data was transformed Gaussian. We implemented this model by setting $\mathbf{H}_{\cdot n}=\phi(Y_{\cdot n}+0.5)$ and proceeding with the uncollapse step of MultiAddGPs directly; skipping sampling the posterior of the collapsed form. We call this model the Normal Additive GP (NAddGP) model.
We compared posterior estimates from the MultiAddGP and NAddGP models to emphasize the importance of modeling uncertainty due to counting. Figure 1 shows that the MultiAddGP almost perfectly recovered the true decomposition whereas the NAddGP substantially underestimated the amplitude of the periodic component and long-term trend.
Figure 1: MultiAddGPs successfully decompose simulated microbiome time-series. The NAddGP model is identical to the MultiAddGP model but ignores uncertainty due to counting by modeling the observed data as transformed Gaussian. Panels A, B, and C represent individual decomposed components associated with each covariate. Panel D illustrates the cumulative effect of all components.
Figure 2 shows these findings generalize as $N$ and $D$ increase. As the posterior of these high-dimensional models cannot be easily visualized, we quantified model performance based on the coverage of posterior 95% intervals with respect to the true function decomposition. Since both MultiAddGP and NAddGP are Bayesian models we do not expect that these intervals will cover the truth with 95% probability. As a result, we focus on the ratio of coverage between the MultiAddGP and the NAddGP. Positive values of this coverage ratio indicate that the MultiAddGP model covers the truth more often than the NAddGP model. In all simulations, at all sample sizes $N$ and number of taxa $D$, the MultiAddGP models covered the truth more frequently than the NAddGP models.
Figure 2: At all tested dimensions $N$ and $D$, posterior intervals from MultiAddGPs cover the truth more frequently than NAddGPs. The first row illustrates how data sparsity varied with dimensions $N$ and $D$ in our simulation studies. The second row shows the ratio between coverage of 95% Credible intervals from MultiAddGPs compared to NAddGPs. Each datapoint represents the mean over three simulations. The ratio is always positive illustrating MultiAddGPs cover the truth substantially more than NAddGPs. Coverage ratios for each of the decomposed components Batch, Hourly, and Daily and the cumulative function $F$ are shown.
5 Analysis of Artificial Human Guts
We used MultiAddGPs to reanalyze on a previously published longitudinal study of four artificial gut models (Silverman et al., 2018). This study consists of over 500 samples from $4$ artificial gut vessels irregularly spaced through time. The key feature of this study was the starvation of Vessels 1 and 2 that occurred over days 11 to 13 while Vessels 3 and 4 received no intervention and therefore act as controls. We have previously modeled these data using Multinomial Logistic-Normal models (Silverman et al., 2022, 2018) and demonstrated that microbiota demonstrate a prolonged recovery phase after feeding resumes. However, prior models were limited and could not separate long-term trend from the effect of starvation. We use a MultiAddGP model to overcome this limitation and provide novel insights into these microbial communities.
Data preprocessing followed the analysis of Silverman et al. (2018) and resulted in 10 bacterial families and 537 samples. We modeled these data as consisting of two overlapping temporal processes, a vessel-specific long-term trend $f^{(\text{base},v)}$ and a vessel-specific function representing the effect of starvation $f^{(\text{disrupt},v)}$. We used the following MultiAddGP model:
$\mathbf{Y}_{\cdot n}$ $\sim\text{Multinomial}(\boldsymbol{\Pi}_{\cdot n})$
$\boldsymbol{\Pi}_{\cdot n}$ $=\text{ALR}_{D}^{-1}(\mathbf{H}_{\cdot n})$
$\mathbf{H}_{\cdot n}$ $\sim N(\mathbf{F}_{\cdot n},\boldsymbol{\Sigma})$
$\mathbf{F}$ $=\sum_{v=1}^{4}(\mathbf{f}^{(\text{base},v)}(t_{n})+I[v\in\{1,2\}]\mathbf{f}^{(\text{disrupt},v)}(t))$
$\mathbf{f}^{(\text{base},v)}$ $\sim\text{GP}(\boldsymbol{\Theta}^{(\text{base},v)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(v)}\odot\boldsymbol{\Gamma}^{(\text{base})}),v\in\{1,2,3,4\}$
$\mathbf{f}^{(\text{disrupt},v)}$ $\sim\text{GP}(\boldsymbol{\Theta}^{(\text{disrupt},v)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(v)}\odot\boldsymbol{\Gamma}^{(\text{disrupt})}),v\in\{1,2\}$
$\boldsymbol{\Sigma}$ $\sim\text{InvWishart}(\boldsymbol{\Xi},\zeta)$
where $\boldsymbol{\Xi}$ was chosen to reflect a prior belief that more evolutionary similar taxa would behave more similarly (Silverman et al., 2017), $\zeta=D+5$ reflects weak prior knowledge in this assumption, and $\odot$ represents elementwise multiplication of kernels. The kernel $\boldsymbol{\Gamma}^{(v)}$ was block diagonal with all non-zero elements equal to 1: it was used to model conditional independence between the vessels. $\boldsymbol{\Gamma}^{(\text{base})}$ was modeled as a squared exponential kernel to capture long-term non-linear trends. In contrast, $\boldsymbol{\Gamma}^{(\text{distrup})}$ was modeled using a rational quadratic kernel which was forced to zero before day 11 to reflect our knowledge that starvation started on day 11. Hyperparameters $\Omega=\{\sigma_{\text{base}},\rho_{\text{base}},\sigma_{\text{disrupt}},\rho_{\text{disrupt}}\}$ were selected using penalized MML. Full details on penalization, kernel, and mean functions are provided in Supplementary Section 5.
Posterior estimates for the cumulative function $F$ (long-term trend plus starvation effect) are shown in Figure 3. Posterior estimates for the decomposed starvation effect are shown in Figure 4. Consistent with prior reports, the Rikenellaceae family demonstrates a substantial decrease during the initial starvation period followed by a slow recovery to baseline (Silverman et al., 2018; Sun and Zhou, 2024). This finding is notable given that decreases in Rikenellaceae are also observed in obese individuals (Okeke, Roland and Mullin, 2014; Peters et al., 2018; Tavella et al., 2021). As obesity and starvation seem like opposite conditions, decreased Rikenellaceae relative abundance in both conditions suggest that host factors not present in these artificial gut systems but present in in vivo obesity studies (e.g., the host immune system) may be driving the decreases in obesity.
Our MultiAddGPs also reveal new features of this data and suggest novel biology. For example, our model finds strong evidence that the Rikenellaceae relative abundance in the community decreases drastically during the initial starvation event followed by a later over-correction where their abundance ultimately increases before trending back to pre-starvation levels. These types of over-corrections are predicted by ecological predator-prey models which can lead to damped oscillations in response to exogenous stimuli (Samuelson, 1971). Still, to best of our knowledge, this is the first study to demonstrate phenomena consistent with that theory within in vitro systems.
Figure 3: MultiAddGP Posterior for the Smoothed State $\mathbf{F}$ of an Artificial Gut Study. As described in the main text, $\mathbf{F}$ was modeled as the additive combination of a slowly-varying temporal trend and a more variable yet short-lived starvation effect in Vessels 1 and 2 which started on Day 11. Posterior mean and 95% credible intervals of $\mathbf{F}$ are shown for each of the four artificial gut vessels. We plot these results with respect to the Centered Log-Ratio (CLR) Coordinates of each bacterial family. The starvation event, which occurred between Days 11 and 13 in Vessels 1 and 2, is highlighted in Gray.
Figure 4: MultiAddGPs Posterior for the Starvation effect ($\mathbf{f}^{(disrupt)}$) in Artificial Gut Study. Posterior mean and 95% credible intervals are shown for all taxa.
6 Discussion
Motivated by applied statistical challenges in the analysis of microbiota data, we have introduced a class of Bayesian Multinomial Logistic-Normal additive linear and non-linear regression models (MultiAddGPs). Using recent theory on Marginally Latent Matrix-t Processes (MLTPs), we have developed efficient inference for this class of models. Beyond MultiAddGPs we also introduced a general framework for hyperparameter selection in MLTPs via Maximum Marginal Likelihood (MML) and penalized MML using the Laplace Approximation within the Collapse-Uncollapse (CU) sampler. These models and methods have been made widely available as part of the fido R package on CRAN. Through both simulated and real data analyses, we have shown that this class of models provides a practical, flexible, and rigorous approach to modeling microbiota data. Still, key challenges and future directions remain.
This article develops an efficient approach to hyperparameter selection via approximate MML and penalized MML estimation. An advantage of our approach is its flexibility: researchers can optimize hyperparameters within a kernel function or even within the log-ratio transformation. In practice, we have found Bayesian Optimization to be a practical approach to maximizing the approximate marginal likelihood with small to moderate numbers of hyperparameters (e.g., ¡10). Further research and improvement are likely needed when selecting larger numbers of hyperparameters simultaneously or when the optimization landscape is particularly challenging.
Besides the challenge of calculating MML and penalized MML estimates, alternative methods of selecting hyperparameters may be needed. In particular, some researchers may require fully Bayesian solutions that can quantify uncertainty in model hyperparameters. This is the notable advantage of Cheng et al. (2019) approach. While their model ignores counting variation, and assumes the data is transformed Gaussian, their method provides fully Bayesian inference of hyperparameters via the slice sampler of Murray and Adams (2010). We expect that the same slice-sampling approach could be used for MultiAddGP, replacing the Metropolis Hastings Steps used in that method with the CU sampler with marginal Laplace approximation.
Finally, readers should be aware of the limits of the Laplace Approximation in the CU sampler when inferring MLN-MLTPs. In Silverman et al. (2022), we proved that the error rate of this approximation is of order $O_{p}([D-1]\sum_{n=1}^{N}[\sum_{d=1}^{D}Y_{dn}]^{-1})$. This result implies that the approximation works well when the average number of counts per sample (e.g., the sequencing depth) is high. This is the case with modern sequencing technologies. In contrast, we expect the approximation will be poor when there are few counts per sample, for example, when observations are categorical ($\sum_{d=1}^{D}Y_{dn}=1$). In the latter case, other approaches are needed.
{acks}
[Acknowledgments] The authors thank Rachel Silverman and Hangzhi Guo for their helpful comments. {funding} TC, MPN, and JDS were supported in part by NIH Grant 1 R01 GM148972-01.
[…]

### v2

@@
-Efficient Bayesian Additive Regression Models For Microbiome Studies
+Scalable Bayesian Semiparametric Additive Regression Models For Microbiome Studies
 Abstract
 Statistical analysis of microbiome data is challenging.
-Bayesian multinomial logistic-normal (MLN) models have gained popularity due to their ability to account for the count compositional nature of these data.
-However, these models are often computationally intractable to infer.
-Recently, we developed a computationally efficient and accurate approach to inferring MLN models with a Marginally Latent Matrix-T Process (MLTP) form: MLN-MLTPs.
-Our approach is based on a novel sampler with a marginal Laplace approximation – called the Collapse-Uncollapse (CU) sampler.
-However, existing work with MLTPs has been limited to linear models or models of a single non-linear process.
-Moreover, existing methods lack an efficient means of estimating model hyperparameters.
-This article addresses both deficiencies.
-We introduce a new class of MLN Additive Gaussian Process models (MultiAddGPs) for deconvolution of overlapping linear and non-linear processes.
-We show that MultiAddGPs are examples of MLN-MLTPs and derive an efficient CU sampler for this model class.
-Moreover, we derive efficient Maximum Marginal Likelihood estimation for hyperparameters in MLTP models by taking advantage of Laplace approximation in the CU sampler.
-We demonstrate our approach using simulated and real data studies.
-Our models produce novel biological insights from a previously published artificial gut study.
-Additive Regression,
-Multinomial Logistic-Normal Models,
-Bayesian Inference,
-Microbiome Data,
-keywords:
-[…]
-, and
+Bayesian multinomial logistic-normal (MLN) models have gained popularity due to their ability to account for the count compositional nature of these data, but existing approaches are either computationally intractable or restricted to purely parametric or non-parametric methods, which limit their flexibility and scalability.
+In this work, we introduce MultiAddGPs, a novel semi-parametric framework that integrates additive Gaussian Process (GP) regression within a Bayesian MLN model to disentangle linear and non-linear covariate effects, including non-stationary dynamics.
+Our approach builds on the computationally efficient Collapse-Uncollapse (CU) sampler and additive GP regression, introducing a novel back-sampling algorithm and marginal likelihood approximation for efficient inference and hyperparameter estimation.
+Our models are over 240,000 times faster than alternatives while simultaneously producing more accurate posterior estimates.
+Additionally, we incorporate non-stationary kernel functions designed to model treatment interventions and disease effects.
+We demonstrate our approach using simulated and real data studies and produce novel biological insights from a previously published human gut microbiome study.
+Our methods are publicly available as part of the fido software package on CRAN  [^1] A preliminary version of this work was presented at the NeurIPS 2024 Workshop on “Bayesian Decision-making and Uncertainty,” but was not published in the workshop proceedings..
+Keywords: Additive Regression, Multinomial Logistic-Normal Models, Bayesian Inference, Microbiome Data, Gaussian Process
 1 Introduction
-Dysregulation of human-, animal-, and even plant-associated microbial communities (microbiota) are known to cause disease (Ballen et al., 2016; Frati et al., 2018; Holleran et al., 2018; Sharon et al., 2019; Gao et al., 2021).
-In humans, alterations of microbiota play a causal role in obesity (Tilg et al., 2011; Ley, 2010), inflammatory bowel disease (Honda and Littman, 2012; Glassner, Abraham and Quigley, 2020; Kostic, Xavier and Gevers, 2014), and even cancer (Schwabe and Jobin, 2013; Helmink et al., 2019).
+Dysregulation of human-, animal-, and even plant-associated microbial communities (microbiota) are known to cause disease (Ballen et al.
+2016, Frati et al.
+2018, Holleran et al.
+2018, Sharon et al.
+2019, Gao et al.
+2021).
+In humans, alterations of microbiota play a causal role in obesity (Tilg et al.
+2011, Ley 2010), inflammatory bowel disease (Honda & Littman 2012, Glassner et al.
+2020, Kostic et al.
+2014), and even cancer (Schwabe & Jobin 2013, Helmink et al.
+2019).
 As a result, many researchers study how dietary, host physiologic, and environmental factors influence the relative abundance of different bacterial taxa in microbiota.
-These factors can have linear or non-linear effects on community structure (Cheng et al., 2019; Sankaran and Jeganathan, 2024; Schwager et al., 2017).
-Even within closed in vitro systems, microbiota can demonstrate non-linear temporal variation (Silverman et al., 2018).
-Overall, flexible statistical methods are needed to disentangle linear and non-linear effects on microbiota.
-Beyond the biological complexity of microbiota, limitations of the measurement process further complicate analyses.
+These factors can have linear or non-linear effects on community structure (Cheng et al.
+2019, Sankaran & Jeganathan 2024, Schwager et al.
+2017).
+Even within closed in vitro systems, microbiota can demonstrate non-linear temporal variation (Silverman et al.
+2018).
+While additive regression models have been proposed to model these complexities, existing implementations are too computationally intensive for widespread use and ignore the count compositional nature of these data (Timonen et al.
+2021, Cheng et al.
+2019).
+Overall, interpretable and flexible statistical methods are needed to disentangle and infer linear and non-linear effects on microbiota.
+Limitations of the sequence counting measurement process complicate the analysis of microbiome data.
 These data are typically represented as a $D\times N$ count table $Y$ with elements $Y_{dn}$ denoting the number of DNA molecules from taxon $d$ observed (sequenced) in sample $n$.
@@
 Instead, the counts are the results of a random sample of that pool of microbes.
-Due to limitations of the measurement process, the size of that sample (the sequencing depth; $\sum_{d=1}^{D}Y_{dn}$) is typically arbitrary and unrelated to the total microbial load in the system (Vandeputte et al., 2017).
-As a result, many authors call these data compositional, reflecting the idea that the data only provide information about the relative abundances of the different taxa within each community (Gloor et al., 2017; McGregor et al., 2023; Mao and Ma, 2022).
-However unlike traditional compositional data (e.g., continuous simplex valued data), these data are zero-laden counts.
-The presence of counting noise and zeros limits standard approaches to compositional data analysis, which typically involve analyses of log-ratio transformed data (Kuczynski et al., 2012; Silverman et al., 2018, 2017, 2020; Kaul et al., 2017; Li, 2015; Gloor et al., 2016).
-Many researchers have turned to Bayesian Multinomial Logistic Normal (MLN) models to address the measurement process’s challenges (Silverman et al., 2021; Äijö, Müller and Bonneau, 2018; Grantham et al., 2020; Silverman et al., 2018, 2022).
-The multinomial is used to model uncertainty due to random counting, while the logistic normal captures the extra-multinomial variability typically seen in these data (Silverman et al., 2022).
-More conceptually, Bayesian MLN models allow researchers to model community composition as a latent simplicial vector that is informed by the observed count data.
-Moreover, unlike the more well-known Dirichlet distribution, the logistic-normal has a rich covariance structure which allows modeling both positive and negative covariation between taxa (Aitchison and Shen, 1980; Silverman et al., 2018).
-The logistic-normal is also self-conjugate (as it is multivariate normal under a suitable log-ratio transformation), allowing for a wide variety of models to be built in the latent simplex space.
-However, the multinomial and the logistic-normal are not conjugate, making inference of these models computationally challenging or even intractable.
-Recent advances have made Bayesian MLN models practical for microbiota analyses.
-Early Bayesian MLN models were inferred via Metropolis-within-Gibbs and only scaled to a handful of dimensions (Cargnoni, Müller and West, 1997).
-Pólya-Gamma data augmentation has been used successfully in some instances (Glynn et al., 2019); however, it is still too computationally intensive for widespread use due to the inability to perform blocked Gibbs updates of Pólya-Gamma random variables while maintaining the logistic-normal form (Linderman, Johnson and Adams, 2015).
-Subsequently, researchers successfully used Hamiltonian Monte Carlo (Äijö, Müller and Bonneau, 2018; Grantham et al., 2020; Silverman et al., 2018).
-Yet scalability still limits those methods.
-For example, the sampler used in Silverman et al.
-(2018) took more than four hours to analyze a dataset of approximately one thousand samples, yet only ten taxa.
-More recently, we proved that a wide variety of Bayesian MLN models, including generalized linear models and generalized Gaussian process regression models, share a common marginal form called a Latent Matrix-T Process (LTP) (Silverman et al., 2022).
-We showed that a Laplace approximation to this marginal form was extremely accurate, leading to an efficient and accurate approximate inference procedure called the Collapse-Uncollapse (CU) sampler.
-Our result demonstrated that this approach is often 4-5 orders of magnitude faster than HMC-based methods with minimal error in posterior calculations (Silverman et al., 2022).
-Despite these advances, there remains a dearth of tools for disentangling the effects of multiple measured factors on microbiota.
-Recently, Cheng et al.
-(2019) proposed an additive Gaussian process framework to address this need.
-Yet their approach assumed the data was transformed Gaussian, ignoring count compositional nature of these data.
-Moreover, our prior work with MLTPs was limited to factors that have a linear effect on microbial composition (generalized linear models) or a single factor that had a nonlinear effect (generalized Gaussian process regression models).
-Finally, our prior work did not address the problem of hyperparameter selection, which is often critical in both linear and nonlinear modeling (e.g., selection of kernel parameters).
-This article addresses the limitations of prior methods and develops a flexible, and computationally efficient approach to disentangling both linear and nonlinear effects on microbiota.
-As in Cheng et al.
-(2019), our approach is based on a class of additive Gaussian process regression models.
-Unlike Cheng et al.
-(2019), we do not assume that the data is transformed Gaussian and instead prove that Bayesian MLN Additive Gaussian Process Models are also MLTPs.
-Using those results, we extend the CU sampler to this class of models.
-Moreover, we extend the results of Silverman et al.
-(2022) by developing an efficient maximum marginal likelihood estimator for model hyperparameters.
-Through simulation studies, we validate our methods; we find posterior intervals from our models often cover the truth ten to a hundred times more often than those of transformed Gaussian models.
-Finally, we apply our models to a previously published longitudinal in vitro gut microbiome study which studied the effects of starvation on microbial communities (Silverman et al., 2018).
-Our models are able to provide the first estimates of the impact of starvation over time within that study.
-We organize this article as follows.
-Section 2 provides a review of the Marginally LTP class of Bayesian MLN models, as well as a review of the Collapse-Uncollapse sampler with marginal Laplace approximation.
-Section 3 presents a multinomial logistic-normal generalized additive Gaussian process regression (MultiAddGP) model and proves that this model is part of the Marginally LTP class.
-We also derive a computationally efficient approach to hyperparameter estimation in MLTP models via maximum marginal likelihood.
-Sections 4 and 5 demonstrate our approach through application to both simulated and real microbiome data.
-Finally, we conclude with a discussion in Section 6.
+Due to limitations of the measurement process, the size of that sample (the sequencing depth; $\sum_{d=1}^{D}Y_{dn}$) is typically arbitrary and unrelated to the total microbial load in the system (Vandeputte et al.
+2017).
+As a result, many authors call these data compositional, reflecting the idea that the data only provide information about the relative abundances of the different taxa within each community (Gloor et al.
+2017, McGregor et al.
+2023, Mao & Ma 2022).
+However, unlike traditional compositional data (e.g., continuous simplex valued data), these data are zero-laden counts.
+The presence of counting noise and zeros limits standard approaches to compositional data analysis, which typically involve analyses of log-ratio transformed data (Kuczynski et al.
+2012, Silverman et al.
+2018, 2017, 2020, Kaul et al.
+2017, Li 2015, Gloor et al.
+2016).
+A growing number of researchers are adopting Bayesian Multinomial Logistic Normal (MLN) models to tackle the challenges associated with the measurement process (Silverman et al.
+2021, Äijö et al.
+2018, Grantham et al.
+2020, Silverman et al.
+2018, 2022, Saxena et al.
+2024).
+The multinomial component addresses uncertainty stemming from random counts, while the logistic normal component accounts for the extra-multinomial variability often observed in these datasets (Silverman et al.
+2022).
+Furthermore, the multinomial framework provides a more intuitive representation of the generating process behind the numerous zero counts than traditional univariate zero-inflated models (Silverman et al.
+2020).
+Conceptually, Bayesian MLN models allow researchers to represent community composition as a latent simplicial vector informed by the observed count data.
+In contrast to the more commonly used Dirichlet distribution, the logistic-normal distribution offers a rich covariance structure that facilitates the modeling of both positive and negative covariation among taxa (Aitchison & Shen 1980, Silverman et al.
+2018).
+The logistic-normal distribution is self-conjugate (as it forms a multivariate normal distribution under a suitable log-ratio transformation), permitting the construction of a diverse range of models within the latent simplex space.
+Although Multinomial-logistic normals have historically posed computational challenges that hinder widespread application, recent innovations, such as the Collapse-Uncollapse (CU) Sampler (Silverman et al.
+2022), have addressed this limitation.
+Nevertheless, there is still a need for more flexible approaches that can effectively disentangle linear and non-linear effects as existing methods do not allow for additive linear and non-linear modeling (Silverman et al.
+2022).
+In this study, we introduce MultiAddGPs, a versatile and interpretable semi-parametric framework aimed at addressing key challenges in microbiome data analysis.
+MultiAddGPs are scalable Bayesian MLN models that facilitate additive combinations of both linear and non-linear Gaussian process components.
+Our approach demonstrates exceptional computational efficiency and scalability, outperforming existing alternatives by a factor of over 400 while yielding more accurate posterior estimates.
+Our inference algorithms leverage the CU Sampler from Silverman et al.
+(2022), while simultaneously enhancing those methods.
+We introduce a novel back-sampling algorithm for efficient posterior simulation from multivariate additive Gaussian process models as well as non-stationary kernels for estimation of non-linear intervention effects on complex microbial communities.
+Moreover, we introduce an accurate and efficient method for marginal likelihood estimation that leverages the Laplace approximation within the CU sampler.
+Our marginal likelihood estimation procedure is critical as it allow for efficient hyperparameter estimation for MultiAddGPs as well as the broader class of models that can be sampled by the CU sampler (Silverman et al.
+2022).
+We validate our methods through simulation and real data studies.
+We demonstrate that MultiAddGPs can directly estimate complex, time-lagged effects of interventions on human microbiota.
+In an application involving an ex vivo artificial human gut study, our findings offer the first evidence that microbiota communities display oscillatory-like behavior in response to short-term dietary interventions.
+MultiAddGPs MLTPs Lgpr LongGP
+Model Linear and Nonlinear Linear or Nonlinear Nonlinear Nonlinear
+Flexible Kernels
+Scalable
+Hyperparameter Estimation
+Count Compositional
+Table 1: A comparison of MultiAddGPs and existing alternatives.
+MultiAddGPs (this article) facilitate additive modeling of both linear and non-linear processes.
+In contrast, MLTPs (Silverman et al.
+2022) consist of either a single linear or a single non-linear component.
+The Lgpr (Timonen et al.
+2021) and LongGP (Cheng et al.
+2019) frameworks permit only additive non-linear components.
+Both MultiAddGPs and MLTPs are implemented in the fido software package available on CRAN, offering both pre-defined and user-defined kernels.
+Conversely, the Lgpr and LongGP packages restrict usage to pre-defined kernels.
+MultiAddGPs and MLTPs are computationally efficient due to the CU Sampler, while LongGP and Lgpr are computationally intensive and do not scale effectively to the datasets analyzed in this article.
+Notably, the MLTPs proposed by Silverman et al.
+(2022) lack a mechanism for estimating model hyperparameters, a feature present in the other three methods.
+Furthermore, MultiAddGPs and MLTPs address the count compositional nature of the data using Bayesian MLN models, whereas Lgpr is limited to univariate likelihoods, and LongGP assumes the data follows a Gaussian distribution.
 Notation
@@
 $\boldsymbol{Y}$) since, in practice, we only evaluate these at a finite number of test points.
-2 Review of Marginally Latent Matrix-T Processes and the CU Sampler
-We describe the class of Marginally Latent Matrix-T Processes (MLTPs) by sequentially generalizing from Matrix-T Processes to Latent Matrix-T Processes (LTPs) and finally MLTPs.
-We then describe the subset of Bayesian Multinomial Logistic-Normal MLTPs (MLN-MLTPs) before reviewing the inference of this class of models.
-2.1 Defining Marginally Latent Matrix-T Processes (MLTPs)
-Just as Gaussian processes can be defined based on the marginal properties of the multivariate normal, matrix normal processes and Matrix-T processes can be defined by the marginal properties of the matrix normal and matrix-T distributions (Silverman et al., 2022).
-Matrix-T processes generalize Student-T processes and Gaussian processes (Silverman et al., 2022).
+2 Methods
+2.1 Multinomial Logistic Normal Additive Gaussian Process Models (MultiAddGPs)
+Let $\mathbf{Y}_{\cdot n}$ denote a $D$-vector of observed data, $\mathbf{X}_{\cdot n}$ a $Q_{0}$-vector of covariates to model linearly, and each $\mathbf{Z}^{(k\in\{1,\dots,K\})}_{\cdot n}$ a $Q_{k}$-vector of covariates to be modeled with distinct non-linear functions.
+The MultiAddGPs model have the following form:
+$\mathbf{Y}_{\cdot n}$ $\sim\text{Multinomial}(\boldsymbol{\Pi}_{\cdot n})$ (1)
+$\boldsymbol{\Pi}_{\cdot n}$ $=\phi^{-1}(\mathbf{H}_{\cdot n})$ (2)
+$\mathbf{H}_{\cdot n}$ $\sim\mathcal{N}(\mathbf{F}_{\cdot n},\boldsymbol{\Sigma})$ (3)
+$\mathbf{F}$ $=\mathbf{B}\mathbf{X}+\sum_{k=1}^{K}\mathbf{f}^{(k)}(\mathbf{Z}^{(k)})$ (4)
+with priors $\mathbf{B}\sim\mathcal{N}(\boldsymbol{\Theta}^{(0)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(0)})$, $\mathbf{f}^{(k)}\sim\mathcal{GP}(\boldsymbol{\Theta}^{(k)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(k)})$, and $\boldsymbol{\Sigma}\sim\mathcal{IW}(\boldsymbol{\Xi},\zeta)$.
+Here, the likelihood is $p(\mathbf{Y}_{\cdot 1},\dots,\mathbf{Y}_{\cdot N})\sim\prod_{n=1}^{N}\text{Multinomial}(\boldsymbol{\Pi}_{\cdot n})$, $\boldsymbol{\Sigma}$ is a $(D-1)\times(D-1)$ covariance matrix, and the transform $\phi$ is an invertible log-ratio transform from the $D$-dimensional simplex to $D-1$ dimentional real-space: $\mathbf{H}_{\cdot n}=\phi(\boldsymbol{\Pi}_{\cdot n}\in\mathbb{S}^{D})\in\mathbb{R}^{D-1}$.
+For computational efficiency, we used the following Additive Log-Ratio (ALR) transform, which takes the $D$-th taxa as a reference:
+$\mathbf{H}_{.n}=\phi(\boldsymbol{\Pi}_{.n})=\left\{\log\left(\frac{\pi_{1n}}{\pi_{Dn}}\right),\ldots,\log\left(\frac{\pi_{(D-1)n}}{\pi_{Dn}}\right)\right\}^{T}.$
+There is no loss of generality as posterior samples taken with respect to the $\text{ALR}_{D}$ coordinate system can be transformed into any other log-ratio coordinate system (Pawlowsky-Glahn et al.
+2015, Appendix A.3).
+For context, the $\text{ALR}_{D}$ transform is the inverse of the identified softmax transform.
+For the matrix-normal prior on the linear term, $\boldsymbol{\Theta}^{(0)}$ is the mean matrix and $\boldsymbol{\Gamma}^{(0)}$ is a $Q_{0}\times Q_{0}$ covariance matrix representing covariance in the parameters of the $Q_{0}$ covariates.
+The terms $\boldsymbol{\Theta}^{(k)}$ and $\boldsymbol{\Gamma}^{(k)}$ in the $K$ matrix-normal process priors echo their linear counterparts but are functions (e.g., mean and kernel functions) rather than fixed dimensional matrices.
+Additionally, we permit hyperparameters $\Omega$ in any part of the model, including but not limited to hyperparameters in the kernel functions (e.g., bandwidth parameters) or hyperparameters in the mean functions.
+As we will show through simulated and real data analyses in Sections 3 and 4, this is a very flexible class of models which can be used in a wide range of additive linear and non-linear modeling tasks.
+2.2 Posterior Estimation in MultiAddGPs
+We develop scalable inference for MultiAddGPs using a two-step Collapse-Uncollapse (CU) sampler (Silverman et al.
+2022).
+Any multi-parameter Bayesian model $p(\mathbf{H},\mathbf{\Phi}\mid\mathbf{Y})$ can be sampled in two-step: (1) obtain a sample from the marginal $p(\mathbf{H}\mid\mathbf{Y})$ (the collapse step); (2) conditioned on that sample, obtain a sample from the corresponding conditional $p(\mathbf{\Phi}\mid\mathbf{H},\mathbf{Y})$ (the uncollapse step).
+Yet this approach is often infeasible as these posterior marginals and conditionals are rarely available for non-conjugate models.
+Full details on the CU sampler can be found in Supplementary Section 1.
+Silverman et al.
+(2022) showed that there was a large class of semi-conjugate models called Marginally Latent matrix-T Processes (MLTPs) where, the posterior conditional $p(\mathbf{\Phi}\mid\mathbf{H},\mathbf{Y})$ had a closed form and the posterior marginal $p(\mathbf{H}\mid\mathbf{Y})$ was the posterior of a Latent matrix-T Process (LTP) which could be accurately and efficiently sampled through a Laplace approximation.
+Below we show that MultiAddGPs are MLTPs and derive the parameters of their marginal LTP form.
+These results, are used to form the Laplace approximation to the Collapse step for MultiAddGPs.
+We then derive a novel Back-sampling (BS) for multivariate additive Gaussian process regression models.
+The BS algorithm provides a closed-form solution for the Uncollapse step for MultiAddGPs.
+Finally, we develop efficient hyperparameter inference for MLTPs (including MultiAddGPs) by maximum marginal likelihood estimation and discuss the identification of posterior samples from MultiAddGPs.
+Algorithm 1 The Collapse-Uncollapse (CU) Sampler for MultiAddGPs Models
+Input: Data $\{\mathbf{Y},\mathbf{X},\mathbf{Z}^{(1)},\dots,\mathbf{Z}^{(K)}\}$ [^2] We omit conditioning on covariates $\mathbf{X}$ and $\mathbf{Z}$ in the notation for brevity., and prior parameters $\Delta=\{\mathbf{\Theta}^{(0)},\dots,\mathbf{\Theta}^{(K)},\mathbf{\Gamma}^{(0)},\dots,\mathbf{\Gamma}^{(K)},\boldsymbol{\Xi},\zeta\}$.
+Output: $S$ samples of $\{\mathbf{H},\boldsymbol{\Sigma},\mathbf{F},\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}\}$.
+1:for each sample $s\in\{1,\dots,S\}$ do
+2:     Sample $\mathbf{H}^{[s]}\sim p(\mathbf{H}|\mathbf{Y},\Delta)$ where $p(\mathbf{H},\mathbf{Y},\Delta)$ is an LTP; $\triangleright$ Collapsed Step
+3:     Sample $\{\boldsymbol{\Sigma}^{[s]},\mathbf{F}^{[s]}\}\sim p(\mathbf{F},\boldsymbol{\Sigma}\mid\mathbf{H}^{[s]},\Delta)$; $\triangleright$ Uncollapsed Step (1)
+4:     Sample $\mathbf{B}^{[s]},\mathbf{f}^{(1)[s]},\dots,\mathbf{f}^{(K)[s]}=\text{Back-Sampling}(\mathbf{Y},\mathbf{F}^{[s]},\boldsymbol{\Sigma}^{[s]},\Delta)$; $\triangleright$ Uncollapsed Step (2)
+5:end for
+2.2.1 MultiAddGPs as MLTPs and the Collapse Step
+Just as Gaussian processes can be defined based on the marginal properties of the multivariate normal, matrix normal processes and matrix-t processes can be defined by the marginal properties of the matrix normal and matrix-t distributions (Silverman et al.
+2022).
+Matrix-t processes generalize Student-t processes and Gaussian processes (Silverman et al.
+2022).
 Definition 2.1 (Matrix-T Process).
-A stochastic process $\mathbf{Y}\sim TP(\nu,\mathbf{M},\mathbf{V},\mathbf{A})$ defined on the set $\mathcal{W}=\mathcal{W}^{(1)}\times\mathcal{W}^{(2)}$ is a matrix-T process if $\mathbf{Y}$ evaluated on any two finite subsets $\mathcal{X}^{(1)}\subset\mathcal{W}^{(1)}$ and $\mathcal{X}^{(2)}\subset\mathcal{W}^{(2)}$ is a random matrix $\mathbf{Y}$ of dimension $|\mathcal{X}^{(1)}|\times|\mathcal{X}^{(2)}|$ that follows a matrix-T distribution: $\mathbf{Y}\sim T(\nu,\mathbf{M,V,A})$.
+A stochastic process $\mathbf{Y}\sim TP(\nu,\mathbf{M},\mathbf{V},\mathbf{A})$ defined on the set $\mathcal{W}=\mathcal{W}^{(1)}\times\mathcal{W}^{(2)}$ is a matrix-T process if $\mathbf{Y}$ evaluated on any two finite subsets $\mathcal{X}^{(1)}\subset\mathcal{W}^{(1)}$ and $\mathcal{X}^{(2)}\subset\mathcal{W}^{(2)}$ is a random matrix of dimension $|\mathcal{X}^{(1)}|\times|\mathcal{X}^{(2)}|$ that follows a matrix-T distribution: $\mathbf{Y}\sim T(\nu,\mathbf{M,V,A})$.
 $\nu$ is a scalar value strictly greater than zero.
@@
 $\mathbf{M}_{ij}=\mathbf{M}(x_{i}^{(1)},x_{j}^{(2)})$ is the matrix function representing the mean, and $\mathbf{V}_{ij}=\mathbf{V}(x_{i}^{(1)},x_{j}^{(1)})$ and $\mathbf{A}_{ij}=\mathbf{A}(x_{i}^{(2)},x_{j}^{(2)})$ are kernel functions.
-Latent Matrix-T Processes (LTP) generalize Matrix-T processes. $\mathbf{Y}$ is said to be an LTP if
-$\mathbf{Y}$ $\sim g(\mathbf{\Pi},\lambda)$
+Latent Matrix-T Processes (LTP) generalize Matrix-T processes.
+Definition 2.2 (Latent Matrix-T Processes).
+$\mathbf{Y}$ is said to be an LTP if
+$\mathbf{Y}$ $\sim g(\mathbf{\Pi})$
 $\mathbf{\Pi}$ $=\phi^{-1}(\mathbf{H})$
 $\mathbf{H}$ $\sim TP(\nu,\mathbf{M},\mathbf{V},\mathbf{A}).$
-where $g$ is any distribution depending on parameters $\mathbf{\Pi}$ as well as hyperparameters $\lambda$ and $\phi$ is a known transform.
-LTPs can alternatively be written as a joint model $p(\mathbf{Y},\mathbf{H})$.
-A stochastic process $\mathbf{Y}$ is Marginally LTP (MLTP) if it can be described by a joint distribution $p(\mathbf{Y},\mathbf{H},\mathbf{\Phi})$ with a marginal $p(\mathbf{Y},\mathbf{H})$ that is a LTP.
-Silverman et al.
-(2022) showed that a wide variety of linear, dynamic linear, and non-linear regression models are MLTP.
-In Section 3, we show that our proposed class of generalized additive Gaussian process regression models are MLTP as well.
-2.2 Bayesian Multinomial Logistic Normal MLTPs (MLN-MLTPs)
-Bayesian Multinomial Logistic Normal MLTPs (MLN-MLTPs) are a subtype of MLTPs that are particularly useful for the analysis of microbiome data.
-In MLN-MLTPs, the distribution $g$ is a product multinomial: $p(\mathbf{Y}_{\cdot 1},\dots,\mathbf{Y}_{\cdot N})\sim\prod_{n=1}^{N}\text{Multinomial}(\boldsymbol{\Pi}_{\cdot n})$ and the transform $\phi$ is an invertible log-ratio transform from the $D$-dimensional simplex to $D-1$ dimentional real-space: $\mathbf{H}_{\cdot n}=\phi(\boldsymbol{\Pi}_{\cdot n}\in\mathbb{S}^{D})\in\mathbb{R}^{D-1}$.
-Canonically, we used the following Additive Log-Ratio (ALR) transform which takes the $D$-th taxa as a reference:
-$\mathbf{H}_{.n}=\phi(\boldsymbol{\Pi}_{.n})=\left\{\log\left(\frac{\pi_{1n}}{\pi_{Dn}}\right),\ldots,\log\left(\frac{\pi_{(D-1)n}}{\pi_{Dn}}\right)\right\}^{T}.$ (1)
-We choose this transform for computational efficiency as discussed in Silverman et al.
-(2022).
-There is no loss in generality as posterior samples taken with respect to the $\text{ALR}_{D}$ coordinate system can be transformed into any other log-ratio coordinate system (Pawlowsky-Glahn, Egozcue and Tolosana-Delgado, 2015, Appendix A.3).
-For context, the $\text{ALR}_{D}$ transform is the inverse of the softmax transform.
-2.3 Collapsed-Uncollapsed (CU) Sampler
-The definition of MLTPs is key to efficient inference.
-If a model $p(\mathbf{Y},\mathbf{H},\mathbf{\Phi})$ has a closed-form marginal $p(\mathbf{Y},\mathbf{H})$ that is an LTP, then its closed form conditional $p(\mathbf{\Phi}\mid\mathbf{Y},\mathbf{H})$ likely exists.
-We call the marginal $p(\mathbf{Y},\mathbf{H})$ the collapsed form and $p(\mathbf{\Phi}\mid\mathbf{Y},\mathbf{H})$ the uncollapsed form.
-The posterior of an MLTP factors as
-$p(\mathbf{H},\Phi\mid\mathbf{Y})=p(\mathbf{\Phi}\mid\mathbf{H},\mathbf{Y})p(\mathbf{H}\mid\mathbf{Y})$
-with the uncollapsed form as the first term and the posterior of the collapsed form as the second.
-As the collapsed form is rarely conjugate, techniques such as MCMC can be used to obtain samples from it’s posterior.
-Then, conditioned on those samples, the uncollapsed form can be used to obtain samples from the joint posterior.
-Especially when $\boldsymbol{\Phi}$ is high-dimensional, this Collapse-Uncollapse sampler can be much more efficient than common alternatives (Silverman et al., 2022).
-Still, the most substantial enhancements occur when approximations to the collapsed form are considered.
-We have developed a Laplace approximation for the collapsed form of MLTPs:
-$p(vec(\mathbf{H})\mid\mathbf{Y})\approx N(vec(\hat{\mathbf{H}}),\nabla^{-2}[vec(\hat{\mathbf{H}})])$
-where $\hat{\mathbf{H}}$ denotes the Maximum A Posteriori (MAP) estimate of the collapsed form and $\nabla^{-2}[vec(\hat{\mathbf{H}})]$ denotes the inverse Hessian of the collapsed form evaluated at the MAP estimate.
-To facilitate this approximation, we derived analytical results for the gradient and Hessian of the collapsed form (Silverman et al., 2022).
-Focusing on applications to MLN-MLTPs, we proved error bounds on the Laplace approximation and provided simulation and real analyses showing that the approximation was extremely accurate in the context of microbiome data analysis.
-Beyond the accuracy of posterior calculations, we showed that this CU sampler with the Laplace approximation (simply referred to as the CU sampler in the following text) was often 4-5 orders of magnitude faster than MCMC and 1-2 orders of magnitude faster than black-box variational inference while also being more accurate than the later.
-The CU sampler for MLN-MLTPs, along with uncollapse samplers for linear and non-linear regression models, is publicly available on CRAN as part of the fido software package.
-3 Methods
-3.1 Multinomial Logistic Normal Additive Gaussian Process Models (MultiAddGP)
-To facilitate additive linear and non-linear modeling using Bayesian MLNs, this article introduces the class of Multinomial Logistic Normal Additive Gaussian Process Models (MultiAddGPs).
-Let $\mathbf{Y}_{\cdot n}$ denote a $D$-vector of observed data, $\mathbf{X}_{\cdot n}$ a $Q_{0}$-vector of covariates to model linearly, and each $\mathbf{Z}^{(k\in\{1,\dots,K\})}_{\cdot n}$ a $Q_{k}$-vector of covariates to be modeled with distinct non-linear functions.
-MultiAddGP models have the following form:
-$\mathbf{Y}_{\cdot n}$ $\sim\text{Multinomial}(\boldsymbol{\Pi}_{\cdot n})$ (2)
-$\boldsymbol{\Pi}_{\cdot n}$ $=\phi^{-1}(\mathbf{H}_{\cdot n})$ (3)
-$\mathbf{H}_{\cdot n}$ $\sim N(\mathbf{F}_{\cdot n},\boldsymbol{\Sigma})$ (4)
-$\mathbf{F}$ $=\mathbf{B}\mathbf{X}+\sum_{k=1}^{K}\mathbf{f}^{(k)}(\mathbf{Z}^{(k)})$ (5)
-with priors $\mathbf{B}\sim N(\boldsymbol{\Theta}^{(0)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(0)})$, $\mathbf{f}^{(k)}\sim\text{GP}(\boldsymbol{\Theta}^{(k)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(k)})$, and $\boldsymbol{\Sigma}\sim\text{InvWishart}(\boldsymbol{\Xi},\zeta)$.
-As in prior sections, $\phi$ denotes any log-ratio transform from $\mathbb{S}^{D}$ to $\mathbb{R}^{D-1}$.
-$\boldsymbol{\Sigma}$ is a $(D-1)\times(D-1)$ covariance matrix.
-For the matrix-normal prior on the linear term, $\boldsymbol{\Theta}^{(0)}$ is the mean matrix and $\boldsymbol{\Gamma}^{(0)}$ is a $Q_{0}\times Q_{0}$ covariance matrix representing covariance in the parameters of the $Q_{0}$ covariates.
-The terms $\boldsymbol{\Theta}^{(k)}$ and $\boldsymbol{\Gamma}^{(k)}$ in the $K$ matrix-normal process priors echo their linear counterparts but are functions (e.g., mean and kernel functions) rather than fixed dimensional matrices.
-Additionally, we permit hyperparameters $\Omega$ in any part of the model including but not limited to hyperparameters in the kernel functions (e.g., bandwidth parameters) or hyperparameters in the mean functions.
-As we will show through simulated and real data analyses in Sections 4 and 5, this is a very flexible form of model which can be used in a wide range of additive linear and non-linear modeling tasks.
-3.2 Posterior Estimation in MultiAddGPs
-We use MLTP theory to sample from the posterior of MultiAddGPs: $p(\mathbf{H},\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)},\boldsymbol{\Sigma}\mid\mathbf{Y})$.
-To connect to that theory we first prove that MultiAddGPs are MLN-MLTP’s with $\boldsymbol{\Phi}=\{\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)},\boldsymbol{\Sigma}\}$.
-We derive the collapsed form and then derive an efficient algorithm for sampling from the uncollapsed form.
-First however, we must clarify the definition of $\mathbf{F},\mathbf{f}_{1},\dots,\mathbf{f}_{K}$.
-Up to this point, we have not distinguished between the set of points $n\in\{1,\dots,N\}$ at which we have observed data $\mathbf{Y}$ and the potentially different set $n^{*}\in\{1,\dots,N^{*}\}$ at which we want to evaluate the functions $\mathbf{F},\mathbf{f}_{1},\dots,\mathbf{f}_{K}$.
-In what follows, we use the symbols $\mathbf{F},\mathbf{f}_{1},\dots,\text{ and }\mathbf{f}_{K}$ to denote the evaluation of corresponding infinite-dimensional functions at the set of evaluation points $\{1,\dots,N^{*}\}$, i.e., they are each $D\times N^{*}$-dimensional random matrices.
-In contrast, all other random matrices (e.g., $\mathbf{H}$) represent their corresponding infinite-dimensional analogues evaluated at the set of observed points ($\{1,\dots,N\}$).
-In Supplementary Section 1 we prove that MultiAddGPs are MLN-MLTPs with the following parameters defining the collapsed form:
-$\mathbf{M}$ $=\boldsymbol{\Theta}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Theta}^{(k)}(\mathbf{Z}^{(k)})$
-$\mathbf{V}$ $=\boldsymbol{\Xi}$
-$\mathbf{A}$ $=\mathbf{X}^{T}\boldsymbol{\Gamma}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}(\mathbf{Z}^{(k)})+\mathbf{I}$
-$\mathbf{\nu}$ $=\mathbf{\zeta}$
-[…]
-Sampling from the uncollapsed form starts by obtaining samples from $p(\mathbf{F},\boldsymbol{\Sigma}\mid\mathbf{H})$.
-Conditioning on $\mathbf{H}$ and marginalizing over $\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}$ in the MultiAddGP model results in a Bayesian matrix-normal process model with likelihood $\mathbf{H}_{\cdot n}\sim N(\mathbf{F}_{\cdot n},\boldsymbol{\Sigma})$ and priors:
-$\mathbf{F}$ $\sim\text{GP}\left(\boldsymbol{\Theta}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Theta}^{(k)}(\mathbf{Z}^{(k)}),\;\boldsymbol{\Sigma},\;\mathbf{X}^{T}\boldsymbol{\Gamma}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}(\mathbf{Z}^{(k)})\right)$
-$\boldsymbol{\Sigma}$ $\sim\text{InvWishart}(\boldsymbol{\Xi},\zeta).$
-This is the same model discussed in Silverman et al. (2022): samples from $p(\mathbf{F},\boldsymbol{\Sigma}\mid\mathbf{H})$ can be obtained via methods described in Appendix C of that article.
-Finally, conditioned on samples of $\mathbf{F}$ and $\boldsymbol{\Sigma}$, we obtain samples of each $\mathbf{f}^{(k)}$.
-Inspired by the backfitting algorithm used for estimation in generalized additive models (Hastie and Tibshirani, 1990), we developed a backsampling algorithm for this task.
-The backsampling proceeds by iteratively sampling $p(\mathbf{B}\mid\mathbf{F})$, $p(\mathbf{f}_{1}\mid\mathbf{F},\mathbf{B})$, $p(\mathbf{f}_{2}\mid\mathbf{F},\mathbf{B},\mathbf{f}_{1})$, $\dots$, and then $p(\mathbf{f}^{(K)}\mid\mathbf{F},\mathbf{B},\mathbf{f}_{1},\dots,\mathbf{f}^{(K-1)})$.
-For brevity, we leave a description of this algorithm to Supplementary Sections 2 and 3.
+where $g$ is any distribution depending on parameters $\mathbf{\Pi}$ and $\phi$ is a known transform.
+A stochastic process $\mathbf{Y}$ is Marginally LTP (MLTP) if there exists a joint distribution $p(\mathbf{Y},\mathbf{H},\mathbf{\Phi})$ such that the marginal $p(\mathbf{Y},\mathbf{H})$ is a LTP.
+In Supplementary Section 2, we prove the following theorem, which establishes that MultiAddGPs are MLTPs.
+Theorem 2.3.
+MultiAddGPs are MLTPs with parameters $\boldsymbol{\Phi}=\{\mathbf{F},\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)},\boldsymbol{\Sigma}\}$.
+The collapsed form $p(\mathbf{Y},\mathbf{H})$ is an LTP:
+$\mathbf{Y}_{\cdot n}$ $\sim\text{Multinomial}(\mathbf{\Pi}_{\cdot n})$
+$\mathbf{\Pi}_{\cdot n}$ $=\phi^{-1}(\mathbf{H}_{\cdot n})$
+$\mathbf{H}$ $\sim TP(\nu,\mathbf{M},\mathbf{V},\mathbf{A})$
+where $\phi$ is an invertible log-ratio transformation, $\mathbf{M}=\boldsymbol{\Theta}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Theta}^{(k)}(\mathbf{Z}^{(k)})$, $\mathbf{V}=\boldsymbol{\Xi}$, and $\mathbf{A}=\mathbf{X}^{T}\boldsymbol{\Gamma}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}(\mathbf{Z}^{(k)})+\mathbf{I}_{N},\nu=\zeta$
+In the collapsed step of the CU sampler, we approximate $p(\mathbf{H}\mid\mathbf{Y})$ using a Laplace approximation: $q(\mathbf{H}|\mathbf{Y})=\mathcal{N}\left(\text{vec}(\hat{\mathbf{H}}),\nabla^{-2}(\text{vec}(\hat{\mathbf{H}}))\right)$, where $\hat{\mathbf{H}}$ is the maximum a posteriori (MAP) estimate of $p(\mathbf{H}\mid\mathbf{Y})$, and $\nabla^{-2}(\text{vec}(\hat{\mathbf{H}}))$ is the inverse Hessian of $\log p(\mathbf{H}\mid\mathbf{Y})$ evaluated at $\hat{\mathbf{H}}$.
+We obtain efficient MAP estimates $\hat{\mathbf{H}}$ by applying the L-BFGS optimizer to the closed-form gradients $\nabla\bigl{(}\mathrm{vec}(\hat{\mathbf{H}})\bigr{)}$ derived in Supplementary Section 3.
+Building on these results, we then compute the Hessian $\nabla^{-2}\bigl{(}\mathrm{vec}(\hat{\mathbf{H}})\bigr{)}$ in closed form (also see Supplementary Section 3).
+This approximation is computationally efficient and extremely accurate.
+The error in this approximation is bounded in probability by $O_{p}([D-1]\sum_{n=1}^{N}\lambda_{n}^{-1})$ where $\lambda_{n}=\sum_{d=1}^{D}\mathbf{Y}_{dn}$ (See Proposition 11 of Silverman et al.
+(2022)).
+This bound explains the high accuracy of the approximation in sequence count applications, such as microbiome data analysis, where the sequencing depth $\lambda_{n}$ is typically large.
+2.2.2 Back-sampling and the Uncollapse Step for MultiAddGPs
+Each posterior sample from $p(\mathbf{H},\mid\mathbf{Y})$ is then conditioned upon and used to sample from $p(\mathbf{\Phi}\mid\mathbf{H},\mathbf{Y})$ where $\boldsymbol{\Phi}=\{\mathbf{F},\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)},\boldsymbol{\Sigma}\}$.
+Since $\mathbf{\Phi}$ is conditionally independent of $\mathbf{Y}$ given $\mathbf{H}$ ($\mathbf{\Phi}\perp\mathbf{Y}\mid\mathbf{H}$), the uncollapse step simplifies to deriving an efficient sampler for $p(\mathbf{\Phi}\mid\mathbf{H})$.
+Further, using the fact that $\{\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}\}\perp\mathbf{H}\mid\{\mathbf{F},\boldsymbol{\Sigma}\}$ we can break the uncollapse step into two simpler steps, with the second conditioned on the results of the first: (1) sample $p(\boldsymbol{\Sigma},\mathbf{F}\mid\mathbf{H})$ and (2) sample $p(\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}\mid\mathbf{F},\boldsymbol{\Sigma})$.
+Before proceeding, we must clarify the distinction between the set of observed data points $n\in\{1,\dots,N\}$, where the response $\mathbf{Y}$ is observed, and the evaluation set $n^{*}\in\{1,\dots,N^{*}\}$, where the functions $\mathbf{F},\mathbf{B}_{0},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}$ are to be inferred.
+In what follows, we use the symbols $\mathbf{F},\mathbf{B}_{0},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}$ to denote the evaluation of corresponding infinite-dimensional functions at the set of evaluation points $\{1,\dots,N^{*}\}$, i.e., they are each $D\times N^{*}$-dimensional random matrices.
+In contrast, all other random matrices (e.g., $\mathbf{H}$) represent their corresponding infinite-dimensional analogues evaluated at the set of observed points $\{1,\dots,N\}$.
+2.2.2.1 Sampling $p(\boldsymbol{\Sigma},\mathbf{F}\mid\mathbf{H})$
+Conditioning on $\mathbf{H}$ and marginalizing over $\mathbf{f}^{(k)},k\in\{1,\dots,K\}$ in the MultiAddGPs model results in a Bayesian matrix-normal process model with likelihood $\mathbf{H}_{\cdot n}\sim\mathcal{N}(\mathbf{F}_{\cdot n},\boldsymbol{\Sigma})$ and priors:
+$\mathbf{F}$ $\sim\mathcal{GP}\left(\boldsymbol{\Theta}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Theta}^{(k)}(\mathbf{Z}^{(k)}),\;\boldsymbol{\Sigma},\;\mathbf{X}^{T}\boldsymbol{\Gamma}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}(\mathbf{Z}^{(k)})\right)$
+$\boldsymbol{\Sigma}$ $\sim\mathcal{IW}(\boldsymbol{\Xi},\zeta).$
+This model is conjugate and posterior samples of $\mathbf{F}$ and $\mathbf{\Sigma}$ can be efficiently obtained as described in Supplementary Section 2.2.
+2.2.2.2 Sampling $p(\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}\mid\mathbf{F},\boldsymbol{\Sigma})$ via Back-sampling
+Inspired by the backfitting algorithm used in generalized additive models (Wood 2017), we introduce a novel Back-sampling (BS) algorithm which iteratively smooths (i.e., decomposes) $\mathbf{F}$ into additive components.
+For each posterior sample, the BS algorithm proceeds by sampling $p(\mathbf{B}\mid\mathbf{F},\boldsymbol{\Sigma})$ and then sequentially sampling $p(\mathbf{f}^{(1)}\mid\mathbf{B},\mathbf{F},\boldsymbol{\Sigma}),\dots,p(\mathbf{f}^{(K-1)}\mid\mathbf{f}^{(1)},...,\mathbf{f}^{(K-2)},\mathbf{B},\mathbf{F},\boldsymbol{\Sigma})$.
+Note the value of the final $\mathbf{f}^{(K)}$ is uniquely determined by the values of $\mathbf{F},\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K-1)}$ and so a sample of $\mathbf{f}^{(K)}$ from the joint distribution can be obtained by calculating $\mathbf{f}^{(K)}=\mathbf{F}-\mathbf{B}\mathbf{X}-\mathbf{f}^{(1)}-\cdots-\mathbf{f}^{(K-1)}$.
+The pseudocode for the BS algorithm is provided in Algorithm 2.
+First we sample from the posterior $p(\mathbf{B}\mid\mathbf{F},\boldsymbol{\Sigma})$.
+To do this we marginalize over $\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}$ and reparameterize as $\mathbf{B}^{\epsilon}:=\mathbf{F}-\sum_{k=1}^{K}\mathbf{\Theta}^{(k)}(\mathbf{Z}^{(k)})$.
+This leads to the following conjugate Bayesian model for $\mathbf{B}$:
+$\mathbf{B}^{\epsilon}$ $\sim\mathcal{MN}(\mathbf{BX},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{*})$
+$\mathbf{B}$ $\sim\mathcal{MN}(\boldsymbol{\Theta}^{(0)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(0)})$
+where $\boldsymbol{\Gamma}^{*}=\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}(\mathbf{Z}^{(k)})$, and $\boldsymbol{\Theta}^{(0)}$, $\boldsymbol{\Gamma}^{(0)}$ denote the prior mean and covariance of $\mathbf{B}$, respectively.
+Using Supplementary Section 4, we have the following closed form for the posterior:
+$\mathbf{B}\mid\mathbf{B}^{\epsilon},\boldsymbol{\Sigma}\sim\mathcal{MN}\Bigg{(}\left[\mathbf{B}^{\epsilon}\boldsymbol{\Gamma}^{-*}\mathbf{X}^{T}+\boldsymbol{\Theta}^{(0)}\boldsymbol{\Gamma}^{-(0)}\right]\left[\mathbf{X}\boldsymbol{\Gamma}^{-*}\mathbf{X}^{T}+\boldsymbol{\Gamma}^{-(0)}\right]^{-1},\boldsymbol{\Sigma},\;\left[\mathbf{X}\boldsymbol{\Gamma}^{-*}\mathbf{X}^{T}+\boldsymbol{\Gamma}^{-(0)}\right]^{-1}\Bigg{)}$ where $\boldsymbol{\Gamma}^{-*}=(\boldsymbol{\Gamma}^{*})^{-1}$ and $\boldsymbol{\Gamma}^{-(0)}=(\boldsymbol{\Gamma}^{(0)})^{-1}$.
+We use a similar trick to sample each nonlinear component $p(\mathbf{f}^{(1)}\mid\mathbf{B},\mathbf{F},\boldsymbol{\Sigma}),\dots,p(\mathbf{f}^{(K-1)}\mid\mathbf{f}^{(1)},...,\mathbf{f}^{(K-2)},\mathbf{B},\mathbf{F},\boldsymbol{\Sigma})$.
+For each $k\in\{1,\dots,K-1\}$, we reparameterize the model as: $\mathbf{f}^{\epsilon_{k}}=\mathbf{F-BX}-\sum_{i=1}^{k-1}\mathbf{f}^{(i)}-\sum_{j=k+1}^{K}\boldsymbol{\Theta}^{(j)}(\mathbf{Z}^{(j)})$.
+Similar to $\mathbf{B}^{\epsilon}$, $\mathbf{f}^{\epsilon_{k}}$ leads to the following conjugate Bayesian model for each $\mathbf{f}^{(k)}$:
+$\mathbf{f}^{\epsilon_{k}}$ $\sim\mathcal{MN}(\mathbf{f}^{(k)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{*})$
+$\mathbf{f}^{(k)}$ $\sim\mathcal{GP}(\boldsymbol{\Theta}^{(k)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(k)})$
+where $\boldsymbol{\Gamma}^{*}=\sum_{j=k+1}^{K}\boldsymbol{\Gamma}^{(j)}$. Again, this can be sampled in closed form using the results of Supplementary Section 4:
+$\mathbf{f}^{(k)}\mid\boldsymbol{\Sigma},\mathbf{f}^{\epsilon_{k}}\sim\mathcal{MN}\Bigg{(}\left[\mathbf{f}^{\epsilon_{k}}\boldsymbol{\Sigma}^{-*}+\boldsymbol{\Theta}^{(k)}\boldsymbol{\Gamma}^{-(k)}\right]\left[\boldsymbol{\Gamma}^{-*}+\boldsymbol{\Gamma}^{-(k)}\right]^{-1},\boldsymbol{\Sigma},\left[\boldsymbol{\Gamma}^{-*}+\boldsymbol{\Gamma}^{-(k)}\right]^{-1}\Bigg{)}$ where $\boldsymbol{\Gamma}^{-*}$ and $\boldsymbol{\Gamma}^{-(k)}$ are short-hand for $(\boldsymbol{\Gamma}^{*})^{-1}$ and $(\boldsymbol{\Gamma}^{(k)})^{-1}$ respectively.
+Finally, the last component $\mathbf{f}^{(K)}$ is uniquely determined as: $\mathbf{f}^{(K)}=\mathbf{F}-\mathbf{BX}-\sum_{k=1}^{K-1}\mathbf{f}^{(k)}$
+Algorithm 2 Back-sampling (BS)
+Input: $\{\mathbf{Y,X,Z}\}$ are data observations; $\{\mathbf{F},\boldsymbol{\Sigma}\}$ are samples from the uncollapsed step (1); $\Delta=\{\mathbf{\Theta}^{(0)},\dots,\mathbf{\Theta}^{(K)},\mathbf{\Gamma}^{(0)},\dots,\mathbf{\Gamma}^{(K)}\}$ is a set of prior inputs.
+Output: $S$ samples of the form $\{\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}\}$
+1:for $s=1$ to $S$ do
+2:     $\mathbf{B}^{\epsilon}=\mathbf{F}-\sum_{k=1}^{K}\boldsymbol{\Theta}^{(k)}(\mathbf{Z}^{(k)})$
+3:     $\boldsymbol{\Gamma}^{*}=\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}$
+4:     $\mathbf{B}|\mathbf{B}^{\epsilon},\boldsymbol{\Sigma}\sim\mathcal{MN}\Bigg{(}\left[\mathbf{B}^{\epsilon}\boldsymbol{\Gamma}^{-*}\mathbf{X}^{T}+\boldsymbol{\Theta}^{(0)}\boldsymbol{\Gamma}^{-(0)}\right]\left[\mathbf{X}\boldsymbol{\Gamma}^{-*}\mathbf{X}^{T}+\boldsymbol{\Gamma}^{-(0)}\right]^{-1},\;\boldsymbol{\Sigma},\;\left[\mathbf{X}\boldsymbol{\Gamma}^{-*}\mathbf{X}^{T}+\boldsymbol{\Gamma}^{-(0)}\right]^{-1}\Bigg{)}$
+5:     for $k=1$ to $K$ do
+6:         $\mathbf{f}^{\epsilon_{k}}=\mathbf{F-BX}-\sum_{i=1}^{k-1}\mathbf{f}^{(i)}-\sum_{j=k+1}^{K}\boldsymbol{\Theta}^{(j)}(\mathbf{Z}^{(j)})$
+7:         $\boldsymbol{\Gamma}^{*}=\sum_{j=k+1}^{K}\boldsymbol{\Gamma}^{(j)}$
+8:         $\mathbf{f}^{(k)}\mid\boldsymbol{\Sigma},\mathbf{f}^{\epsilon_{k}}\sim\mathcal{MN}\left(\left[\mathbf{f}^{\epsilon_{k}}\boldsymbol{\Sigma}^{-*}+\boldsymbol{\Theta}^{(k)}\boldsymbol{\Gamma}^{-(k)}\right]\left[\boldsymbol{\Gamma}^{-*}+\boldsymbol{\Gamma}^{-(k)}\right]^{-1},\;\boldsymbol{\Sigma},\;\left[\boldsymbol{\Gamma}^{-*}+\boldsymbol{\Gamma}^{-(k)}\right]^{-1}\right)$
+9:     end for
+10:     $\mathbf{f}^{(K)}=\mathbf{F}-\mathbf{BX}-\sum_{k=1}^{K-1}\mathbf{f}^{(k)}$
+11:end for
 2.3 Model Identification
@@
 $\tilde{\mathbf{f}}^{(k)}=\mathbf{f}^{(k)}(z^{(k)})-\frac{1}{n}\sum_{n=1}^{N}\mathbf{f}^{(k)}(z_{n}^{(k)}).$
+We find this method is computationally efficient, easy to implement, and provides similar results to alternative methods.
 2.4 Inferring Model Hyperparameters
-Hyperparameter selection remains an outstanding problem in MLTPs (Silverman et al., 2022).
+Hyperparameter selection remains an outstanding problem in MLTPs (Silverman et al.
+2022).
 Here we demonstrate how the same Laplace approximation used in the Collapse step of the CU sampler can be used to approximate marginal likelihoods and thereby estimate hyperparameters via Maximum Marginal Likelihood (MML) or penalized MML estimation.
-Let $\Omega$ denote a set of hyperparameters in the joint MLTP distribution: $p(\mathbf{H},\Phi,\mathbf{Y}\mid\Omega)$.
+Let $\Omega$ denote a set of hyperparameters in the joint MLTPs distribution: $p(\mathbf{H},\Phi,\mathbf{Y}\mid\Omega)$.
 These can be parameters in a kernel, a mean function, or even in the log-ratio transformation $\phi$.
@@
 where $\hat{\mathbf{H}}_{\Omega}$ denotes the MAP estimate for $\mathbf{H}$ evaluated in the observed set $\{1,\dots,N\}$ which depends on the hyperparameters $\Omega$.
-From a computational standpoint, this Laplace approximation to the marginal likelihood is essentially free given that we are already computing the MAP estimate and a factorization of the Hessian matrix as part of the CU sampler (Silverman et al., 2022).
-While the MML estimate already provides a degree of penalization (Wood, 2011), it can be helpful to add an additional penalty term to the marginal likelihood when defining weakly identified MultiAddGPs.
-As we demonstrate in Section 5 and Supplementary Section 5, we often choose this penalty term as the log density of a prior for the hyperparameters.
-In practice, we find that the CU sampler is efficient enough to make optimizing Equation (7) (with or without added penalization) practical.
-While we have found that various optimization routines can be used to perform the optimization, the problem is often non-convex, especially when estimating bandwidth parameters in kernel functions.
-For this reason, we generally recommend Bayesian optimization procedures which can optimize non-convex functions while simultaneously reducing the number of function evaluations compared to more typical optimizers (e.g., the L-BFGS optimizer).
+From a computational standpoint, this Laplace approximation to the marginal likelihood is essentially free given that we are already computing the MAP estimate and a factorization of the Hessian matrix during the CU sampler.
+Beyond MML estimation, it is simple to add additional penalization when defining weakly identified MultiAddGPs.
+In practice, we estimate hyperparameters by maximizing Equation (6), either with or without additional penalization.
+While various optimization routines can be used to perform the optimization, the problem is often non-convex, especially when estimating bandwidth parameters in kernel functions.
+For this reason, we generally recommend Bayesian optimization procedures, which can optimize non-convex functions while simultaneously reducing the number of function evaluations compared to more typical optimizers (e.g., the L-BFGS optimizer).
 We have tested and found Bayesian optimization procedures generally work well when $\Omega$ is of low to moderate dimension (e.g., $\leq 10$).
 When $\Omega$ is higher dimensional, alternative procedures may be needed.
-In this study, Bayesian Optimization was performed using the ParBayesianOptimization package with the acquisition function set to the upper confidence bound, starting with 10 initial points and proceeding through 20 iterations with all other parameters set to their default values.
+In this study, Bayesian Optimization was performed using the ParBayesianOptimization R package with the acquisition function set to the upper confidence bound, starting with 10 initial points and proceeding through 20 iterations with all other parameters set to their default values.
 2.5 Software Availability
-We extend the R package fido (Silverman et al.
-(2022)) to efficiently infer MultiAddGPs models.
+We have made MultiAddGPs available through the R package fido (Silverman & Nixon 2021), where it can be accessed via the basset function.
 The fido package implements the extended CU sampler described above using optimized C++ code.
 MAP estimation of $\hat{H}$ is performed using the L-BFGS optimizer.
-Hyperparameter selection by MML is performed by maximizing the marginal likelihood using Bayesian optimization implemented in the ParBayesianOptimization package.
-All code required to reproduce the results of the this article is available at https://github.com/Silverman-Lab/MultiAddGPs.
+Hyperparameter selection by MML is performed by maximizing the marginal likelihood using Bayesian optimization implemented in the ParBayesianOptimization R package (Wilson 2022).
+All code required to reproduce the results of this article is available at https://github.com/Silverman-Lab/MultiAddGPs.
 3 Simulations
-We simulated a suite of longitudinal studies of microbiota with varying numbers of taxa $D\in\{3,\dots,100\}$ and samples $N\in\{20,\dots,1000\}$.
+To evaluate the performance of the proposed model relative to existing methods, we conduct two simulation studies.
+The first study highlights the importance of accounting for uncertainty introduced by the measurement process, which is ignored by LongGP (Cheng et al.
+2019) .
+The second study compares the performance of MultiAddGPs to Lgpr (Timonen et al.
+2021).
+However, due to the high computational cost of Lgpr, we were unable to scale it to the full simulation setting.
+Instead, we perform a smaller-scale simulation to compare MultiAddGPs and Lgpr.
+Full simulation details are provided in Supplementary Section 5.
+3.1 Simulation 1
+We simulated longitudinal microbiome studies with varying numbers of taxa $D\in\{3,\dots,100\}$ and samples $N\in\{50,\dots,1000\}$.
 We simulated microbial composition influenced by batch effects, daily periodicity (e.g., circadian rhythm; Heddes et al.
 (2022)), and longer-term trends.
-Full simulation details are provided in Supplementary Section 4.
+For this simulation study we did not compare to Lgpr or LongGP for two reasons: (1) it was too computationally intensive and could not scale to the larger simulation sizes; (2) it could not model periodic components as it lacked periodic kernels.
+Instead, we created a new method Normal Additive GPs (NAddGPs) which assumes the data are transformed Gaussian and are therefore reminiscent of LongGP.
+NAddGPs are nearly identical to MultiAddGPs but assume each $\mathbf{H}_{\cdot n}$ is observed and equal to $\mathbf{H}_{\cdot n}=\phi(\mathbf{Y}_{\cdot n}+0.5\cdot\mathbf{1}_{D})$ and therefore proceed directly to the uncollapse step, skipping the collapsed step of our CU sampler.
 In Figure 1 we show a small simulation $D=4$ and $N=600$ for ease of visualization.
 We use $t_{n}$ to denote the time at which sample $n$ was obtained.
-For inference, we specify a MultiAddGP model $\mathbf{F}_{n}=\mathbf{b}_{0}+\mathbf{b}_{1}x^{(\text{batch})}_{n}+\mathbf{f}^{(\text{periodic})}(t_{n})+\mathbf{f}^{(\text{trend})}(t_{n})$ as follows.
+For inference, we specified MultiAddGPs and NAddGPs models as $\mathbf{F}_{n}=\mathbf{b}_{0}+\mathbf{b}_{1}x^{(\text{batch})}_{n}+\mathbf{f}^{(\text{periodic})}(t_{n})+\mathbf{f}^{(\text{trend})}(t_{n})$.
 For covariates, we set $\mathbf{X}_{\cdot n}=[1\;x^{(\text{batch})}_{n}]^{T}$, $\mathbf{Z}^{(\text{periodic})}_{\cdot n}=t_{n}$, and $\mathbf{Z}^{(\text{trend})}_{\cdot n}=t_{n}$.
@@
 Both $\mathbf{f}^{(\text{periodic})}$ and $\mathbf{f}^{(\text{trend})}$ were given matrix-normal process priors with mean function $\boldsymbol{\Theta}^{(k)}=\mathbf{0}$.
-A periodic kernel $K_{\text{period}}(t,t^{\prime})=\sigma_{\text{period}}\exp\left(-\frac{2\sin^{2}\left(\frac{\pi|t-t^{\prime}|}{p}\right)}{\rho_{\text{period}}^{2}}\right)$ was used in the prior for $\mathbf{f}^{(\text{periodic})}$ and a squared exponential $K_{\text{trend}}(t,t^{\prime})=\sigma_{\text{trend}}^{2}\exp(-\frac{(t-t^{\prime})^{2}}{2\rho_{\text{trend}}^{2}})$ was used for $\mathbf{f}^{(\text{trend})}$.
+A periodic kernel $K_{\text{period}}(t,t^{\prime})=\sigma_{\text{period}}\exp\left(-\frac{1}{\rho_{\text{period}}^{2}}2\sin^{2}\left(\frac{\pi|t-t^{\prime}|}{p}\right)\right)$ was used in the prior for $\mathbf{f}^{(\text{periodic})}$ and a squared exponential $K_{\text{trend}}(t,t^{\prime})=\sigma_{\text{trend}}^{2}\exp(-\frac{1}{2\rho_{\text{trend}}^{2}}(t-t^{\prime})^{2})$ was used for $\mathbf{f}^{(\text{trend})}$.
 Hyperparameters $\Omega=\{\sigma_{\text{period}},\rho_{\text{period}},p,\sigma_{\text{trend}},\rho_{\text{trend}}\}$ were selected using MML estimation.
-For comparison, we created a nearly identical model that ignored uncertainty due to counting and assumed the data was transformed Gaussian.
-We implemented this model by setting $\mathbf{H}_{\cdot n}=\phi(Y_{\cdot n}+0.5)$ and proceeding with the uncollapse step of MultiAddGPs directly; skipping sampling the posterior of the collapsed form.
-We call this model the Normal Additive GP (NAddGP) model.
-We compared posterior estimates from the MultiAddGP and NAddGP models to emphasize the importance of modeling uncertainty due to counting.
-Figure 1 shows that the MultiAddGP almost perfectly recovered the true decomposition whereas the NAddGP substantially underestimated the amplitude of the periodic component and long-term trend.
+We compared posterior estimates from the MultiAddGPs and NAddGPs models to emphasize the importance of modeling uncertainty due to counting.
+Figure 1 shows that the MultiAddGPs almost perfectly recovered the true decomposition whereas the NAddGPs substantially underestimated the amplitude of the periodic component and long-term trend.
 Figure 1: MultiAddGPs successfully decompose simulated microbiome time-series.
-The NAddGP model is identical to the MultiAddGP model but ignores uncertainty due to counting by modeling the observed data as transformed Gaussian.
+The NAddGPs model is identical to the MultiAddGPs model but ignores uncertainty due to counting by modeling the observed data as transformed Gaussian.
 Panels A, B, and C represent individual decomposed components associated with each covariate.
 Panel D illustrates the cumulative effect of all components.
+In all panels, MultiAddGPs more accurately recover the true underlying signal compared to NAddGPs.
 Figure 2 shows these findings generalize as $N$ and $D$ increase.
 As the posterior of these high-dimensional models cannot be easily visualized, we quantified model performance based on the coverage of posterior 95% intervals with respect to the true function decomposition.
-Since both MultiAddGP and NAddGP are Bayesian models we do not expect that these intervals will cover the truth with 95% probability.
-As a result, we focus on the ratio of coverage between the MultiAddGP and the NAddGP.
-Positive values of this coverage ratio indicate that the MultiAddGP model covers the truth more often than the NAddGP model.
-In all simulations, at all sample sizes $N$ and number of taxa $D$, the MultiAddGP models covered the truth more frequently than the NAddGP models.
+Since both MultiAddGPs and NAddGPs are Bayesian models, we do not expect that these intervals will cover the truth with 95% probability.
+As a result, we focus on the ratio of coverage between the MultiAddGPs and the NAddGPs.
+Positive values of this coverage ratio indicate that the MultiAddGPs model covers the truth more often than the NAddGPs model.
+In all simulations, at all sample sizes $N$ and number of taxa $D$, the MultiAddGPs models covered the truth far more frequently than the NAddGPs models.
 Figure 2: At all tested dimensions $N$ and $D$, posterior intervals from MultiAddGPs cover the truth more frequently than NAddGPs.
@@
 Each datapoint represents the mean over three simulations.
-The ratio is always positive illustrating MultiAddGPs cover the truth substantially more than NAddGPs.
-Coverage ratios for each of the decomposed components Batch, Hourly, and Daily and the cumulative function $F$ are shown.
-5 Analysis of Artificial Human Guts
-We used MultiAddGPs to reanalyze on a previously published longitudinal study of four artificial gut models (Silverman et al., 2018).
+The ratio is always positive, illustrating that MultiAddGPs cover the truth substantially more than NAddGPs.
+Coverage ratios for each of the decomposed components Batch, Hourly, and Daily and the cumulative function $\mathbf{F}$ are shown.
+3.2 Simulation 2
+To compare with Lgpr, which allows for negative binomial likelihoods, we conducted a smaller simulation with $D=3$ taxa and $N=50$ samples.
+Since Lgpr employs predefined kernel structures and does not incorporate a periodic kernel, we designed the simulation based on an alternative latent function $\mathbf{F}_{n}=\mathbf{b_{0}}+\mathbf{b_{1}}x_{n}^{(\text{batch})}+\mathbf{f}^{(\text{non-stationary})}(t_{n})+\mathbf{f}^{(\text{trend})}(t_{n})$.
+We chose priors for Lgpr that matched the priors used for our MultiAddGPs.
+Lgpr required approximately seven hours to analyze a single taxon on a high-performance computing cluster, resulting in an effective sample size of only 100 (Gelman et al.
+2015).
+In contrast, MultiAddGPs simultaneously analyzed all taxa in about one minute and produced an effective sample size of 2000.
+In summary, our method produces effective samples at over 240,000 times the rate of Lgpr.
+Our methods not only demonstrate increased efficiency, but the posterior estimates obtained using MultiAddGPs also align more closely with the true function decomposition compared to those derived from Lgpr (see Supplementary Figure 3).
+4 Application to Artificial Human Guts
+We used MultiAddGPs to reanalyze a previously published longitudinal study of four artificial gut models (Silverman et al.
+2018).
 This study consists of over 500 samples from $4$ artificial gut vessels irregularly spaced through time.
 The key feature of this study was the starvation of Vessels 1 and 2 that occurred over days 11 to 13 while Vessels 3 and 4 received no intervention and therefore act as controls.
-We have previously modeled these data using Multinomial Logistic-Normal models (Silverman et al., 2022, 2018) and demonstrated that microbiota demonstrate a prolonged recovery phase after feeding resumes.
-However, prior models were limited and could not separate long-term trend from the effect of starvation.
-We use a MultiAddGP model to overcome this limitation and provide novel insights into these microbial communities.
+The study were previously be analyzed by Silverman et al.
+(2022) and Silverman et al.
+(2018), who show that microbiota demonstrate a prolonged recovery phase after feeding resumes.
+However, previous models could not disentangle long-term temporal trends from the specific effects of starvation, leaving the dynamics of the recovery phase poorly understood.
+We use a MultiAddGPs model to overcome this limitation and provide novel insights into these microbial communities.
 Data preprocessing followed the analysis of Silverman et al.
@@
 We modeled these data as consisting of two overlapping temporal processes, a vessel-specific long-term trend $f^{(\text{base},v)}$ and a vessel-specific function representing the effect of starvation $f^{(\text{disrupt},v)}$.
-We used the following MultiAddGP model:
+We used the following MultiAddGPs model:
 $\mathbf{Y}_{\cdot n}$ $\sim\text{Multinomial}(\boldsymbol{\Pi}_{\cdot n})$
@@
 $\boldsymbol{\Sigma}$ $\sim\text{InvWishart}(\boldsymbol{\Xi},\zeta)$
-where $\boldsymbol{\Xi}$ was chosen to reflect a prior belief that more evolutionary similar taxa would behave more similarly (Silverman et al., 2017), $\zeta=D+5$ reflects weak prior knowledge in this assumption, and $\odot$ represents elementwise multiplication of kernels.
+where $\boldsymbol{\Xi}$ was chosen following Silverman et al.
+(2022) to reflect the phylogenetic relationships between taxa.
+$\zeta=D+5$ was chosen to reflect weak prior knowledge of the importance of these relationships.
+Within the kernel functions, $\odot$ represents elementwise multiplication of kernels.
 The kernel $\boldsymbol{\Gamma}^{(v)}$ was block diagonal with all non-zero elements equal to 1: it was used to model conditional independence between the vessels.
 $\boldsymbol{\Gamma}^{(\text{base})}$ was modeled as a squared exponential kernel to capture long-term non-linear trends.
-In contrast, $\boldsymbol{\Gamma}^{(\text{distrup})}$ was modeled using a rational quadratic kernel which was forced to zero before day 11 to reflect our knowledge that starvation started on day 11.
+In contrast, $\boldsymbol{\Gamma}^{(\text{distrup})}$ was modeled using a rational quadratic kernel, which was forced to zero before day 11 to reflect our knowledge that starvation started on day 11.
 Hyperparameters $\Omega=\{\sigma_{\text{base}},\rho_{\text{base}},\sigma_{\text{disrupt}},\rho_{\text{disrupt}}\}$ were selected using penalized MML.
@@
 Posterior estimates for the decomposed starvation effect are shown in Figure 4.
-Consistent with prior reports, the Rikenellaceae family demonstrates a substantial decrease during the initial starvation period followed by a slow recovery to baseline (Silverman et al., 2018; Sun and Zhou, 2024).
-This finding is notable given that decreases in Rikenellaceae are also observed in obese individuals (Okeke, Roland and Mullin, 2014; Peters et al., 2018; Tavella et al., 2021).
-As obesity and starvation seem like opposite conditions, decreased Rikenellaceae relative abundance in both conditions suggest that host factors not present in these artificial gut systems but present in in vivo obesity studies (e.g., the host immune system) may be driving the decreases in obesity.
+Consistent with prior reports, the Rikenellaceae family demonstrates a substantial decrease during the initial starvation period, followed by a slow recovery to baseline (Silverman et al.
+2018, Sun & Zhou 2024).
+This finding is notable given that decreases in Rikenellaceae are also observed in obese individuals (Okeke et al.
+2014, Peters et al.
+2018, Tavella et al.
+2021).
+As obesity and starvation seem like opposite conditions, decreased Rikenellaceae relative abundance in both conditions suggests that host factors not present in these artificial gut systems but present in in vivo obesity studies (e.g., the host immune system) may be driving the decreases in obesity.
 Our MultiAddGPs also reveal new features of this data and suggest novel biology.
-For example, our model finds strong evidence that the Rikenellaceae relative abundance in the community decreases drastically during the initial starvation event followed by a later over-correction where their abundance ultimately increases before trending back to pre-starvation levels.
-These types of over-corrections are predicted by ecological predator-prey models which can lead to damped oscillations in response to exogenous stimuli (Samuelson, 1971).
-Still, to best of our knowledge, this is the first study to demonstrate phenomena consistent with that theory within in vitro systems.
-Figure 3: MultiAddGP Posterior for the Smoothed State $\mathbf{F}$ of an Artificial Gut Study.
+For example, our model finds strong evidence that the Rikenellaceae relative abundance in the community decreases drastically during the initial starvation event, followed by a later over-correction where their abundance ultimately increases before trending back to pre-starvation levels.
+These types of over-corrections are predicted by ecological predator-prey models, which can lead to damped oscillations in response to exogenous stimuli (Samuelson 1971).
+To best of our knowledge, this is the first study to demonstrate phenomena consistent with that theory within in vitro systems.
+Figure 3: MultiAddGPs Posterior for the Smoothed State $\mathbf{F}$ of an Artificial Gut Study.
 As described in the main text, $\mathbf{F}$ was modeled as the additive combination of a slowly-varying temporal trend and a more variable yet short-lived starvation effect in Vessels 1 and 2 which started on Day 11.
@@
 The starvation event, which occurred between Days 11 and 13 in Vessels 1 and 2, is highlighted in Gray.
-Figure 4: MultiAddGPs Posterior for the Starvation effect ($\mathbf{f}^{(disrupt)}$) in Artificial Gut Study. Posterior mean and 95% credible intervals are shown for all taxa.
+Figure 4: MultiAddGPs Posterior for the Starvation effect ($\mathbf{f}^{\text{(disrupt)}}$) in Artificial Gut Study. Posterior mean, 50% and 95% credible intervals are shown for all taxa.
+5 Application to Longitudinal Microbiome Data
+We reanalyzed a human gut microbiome study wherein the study participant collected daily fecal samples over the course of one year (David, Materna, Friedman, Campos-Baptista, Blackburn, Perrotta, Erdman & Alm 2014).
+While the original dataset had samples from two donors, we focus on the gut microbiome data from Subject A, as this subject demonstrated a substantial shift in community composition during a period of prolonged international travel (David, Materna, Friedman, Campos-Baptista, Blackburn, Perrotta, Erdman & Alm 2014).
+While this study has been analyzed by multiple authors (Gibbons et al.
+2017, Duvallet et al.
+2017), we are the first to produce direct estimates of the effect of travel, partitioning those effects from background temporal variation (see Figure 5).
+Data was obtained from European Bioinformatics Institute (EBI) with accession number SAMEG179160 and processed following the methods of David, Materna, Friedman, Campos-Baptista, Blackburn, Perrotta, Erdman & Alm (2014), with the exception that we used DADA2 (Callahan et al.
+2016) to call sequence variants rather than the OTU calling methods used in the original study.
+In total, we analyzed 47 bacterial families over 263 time-points.
+We modeled these data as consisting of two overlapping temporal components: a long-term trend, denoted as $\mathbf{f}^{(\text{trend)}}$, and a non-stationary function representing the effect of travel intervention, denoted as $\mathbf{f^{(\text{travel})}}$.
+The data were analyzed using the following MultiAddGPs model:
+$\mathbf{Y}_{\cdot n}$ $\sim\text{Multinomial}(\boldsymbol{\Pi}_{\cdot n})$
+$\boldsymbol{\Pi}_{\cdot n}$ $=\text{ALR}_{D}^{-1}(\mathbf{H}_{\cdot n})$
+$\mathbf{H}_{\cdot n}$ $\sim N(\mathbf{F}_{\cdot n},\boldsymbol{\Sigma})$
+$\mathbf{F}$ $=\beta_{0}+\mathbf{f}^{(\text{trend})}(t_{n})+\mathbf{f}^{(\text{travel})}(t_{n})$
+$\mathbf{f}^{(\text{trend})}$ $\sim\text{GP}(\boldsymbol{\Theta}^{(\text{trend})},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(\text{trend})})$
+$\mathbf{f}^{(\text{travel})}$ $\sim\text{GP}(\boldsymbol{\Theta}^{(\text{travel},v)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(\text{travel})})$
+$\boldsymbol{\Sigma}$ $\sim\text{InvWishart}(\boldsymbol{\Xi},\zeta)$
+where $\boldsymbol{\Xi}$ was chosen following Silverman et al.
+(2022) to reflect the phylogenetic relationships between taxa.
+$\zeta=D+100$ was chosen to reflect moderately strong prior knowledge of the importance of these relationships.
+A squared exponential kernel ($\mathbf{\Gamma}^{\text{trend}}$) was used to capture long-term nonlinear trends, while a non-stationary kernel with a non-linear warp function ($\mathbf{\Gamma}^{\text{travel}}$) used to capture the shorter-term non-linear effect of travel.
+Hyperparameters $\Omega=\{\sigma_{\text{trend}},\rho_{\text{trend}},\sigma_{\text{travel}},\rho_{\text{travel}}\}$ were estimated using penalized MML.
+Figure 5 shows the contribution of each component and the cumulative effects.
+Consistent with David, Maurice, Carmody, Gootenberg, Button, Wolfe, Ling, Devlin, Varma, Fischbach et al.
+(2014), our findings indicate that while the microbiome remains relatively stable over the long term, travel events induce significant perturbations in community composition.
+Full details of the analysis are provided in Supplementary Section 7.
+Figure 5: Posterior of each component and their cumulative effect for bacterial families Fusobacteriaceae, Monoglobaceae, and Yersiniaceaethe (Longitudinal microbiome data).
+Each row shows contribution of individual components ($\mathbf{f}^{\text{(trend)}}$ and $\mathbf{f}^{\text{(travel)}}$) and the final column presents the cumulative effect ($\mathbf{f}^{\text{(trend)}}+\mathbf{f}^{\text{(travel)}}$).
+Results are reported in terms of Centered Log-Ratio (CLR) coordinates, along with posterior means and corresponding 50% as well as 95% credible intervals.
 6 Discussion
 Motivated by applied statistical challenges in the analysis of microbiota data, we have introduced a class of Bayesian Multinomial Logistic-Normal additive linear and non-linear regression models (MultiAddGPs).
-Using recent theory on Marginally Latent Matrix-t Processes (MLTPs), we have developed efficient inference for this class of models.
-Beyond MultiAddGPs we also introduced a general framework for hyperparameter selection in MLTPs via Maximum Marginal Likelihood (MML) and penalized MML using the Laplace Approximation within the Collapse-Uncollapse (CU) sampler.
-These models and methods have been made widely available as part of the fido R package on CRAN.
-Through both simulated and real data analyses, we have shown that this class of models provides a practical, flexible, and rigorous approach to modeling microbiota data.
+Using recent theory on Marginally Latent Matrix-t Processes (MLTPs), we have developed efficient inference and hyperparameter estimation algorithms for this class of models.
+Additionally, we have incorporated non-stationary kernel functions designed to model treatment interventions and disease effects, allowing capture of complex microbiome dynamics.
+Beyond MultiAddGPs, our hyperparameter estimation techniques can be applied in the much larger class of MLTPs that can be sampled using the CU Sampler.
+Our methods have been made widely available as part of the fido R package, which is published on CRAN.
+While our methods are inspired by challenges in microbiome data, MultiAddGPs could be applied broadly to other settings where count compositional data is collected, including gene expression studies (Fernandes et al.
+2014), macro-ecology (Billheimer et al.
+2001), and even political polling (Topîrceanu & Precup 2020).
 Still, key challenges and future directions remain.
@@
 An advantage of our approach is its flexibility: researchers can optimize hyperparameters within a kernel function or even within the log-ratio transformation.
-In practice, we have found Bayesian Optimization to be a practical approach to maximizing the approximate marginal likelihood with small to moderate numbers of hyperparameters (e.g., ¡10).
+In practice, we have found Bayesian Optimization to be a practical approach to maximizing the approximate marginal likelihood with small to moderate numbers of hyperparameters (e.g., <10).
 Further research and improvement are likely needed when selecting larger numbers of hyperparameters simultaneously or when the optimization landscape is particularly challenging.
@@
 (2019) approach.
-While their model ignores counting variation, and assumes the data is transformed Gaussian, their method provides fully Bayesian inference of hyperparameters via the slice sampler of Murray and Adams (2010).
-We expect that the same slice-sampling approach could be used for MultiAddGP, replacing the Metropolis Hastings Steps used in that method with the CU sampler with marginal Laplace approximation.
-Finally, readers should be aware of the limits of the Laplace Approximation in the CU sampler when inferring MLN-MLTPs.
+While their model ignores counting variation, and assumes the data is transformed Gaussian, their method provides fully Bayesian inference of hyperparameters via the slice sampler of Murray & Adams (2010).
+We expect that the same slice-sampling approach could be used for MultiAddGPs, replacing the Metropolis-Hastings Steps used in that method with the CU sampler with marginal Laplace approximation.
+Finally, readers should be aware of the limits of the Laplace Approximation in the CU sampler when inferring Multinomial Logistic-Normal MLTPs.
 In Silverman et al.
-(2022), we proved that the error rate of this approximation is of order $O_{p}([D-1]\sum_{n=1}^{N}[\sum_{d=1}^{D}Y_{dn}]^{-1})$.
+(2022), we proved that the error rate of this Laplace approximation is of order $O_{p}([D-1]\sum_{n=1}^{N}[\sum_{d=1}^{D}Y_{dn}]^{-1})$.
 This result implies that the approximation works well when the average number of counts per sample (e.g., the sequencing depth) is high.
@@
 In contrast, we expect the approximation will be poor when there are few counts per sample, for example, when observations are categorical ($\sum_{d=1}^{D}Y_{dn}=1$).
-In the latter case, other approaches are needed.
-{acks}
-[Acknowledgments] The authors thank Rachel Silverman and Hangzhi Guo for their helpful comments. {funding} TC, MPN, and JDS were supported in part by NIH Grant 1 R01 GM148972-01.
-[…]
-[…]
+In this scenario, alternative methods, such as the Debiased Multinomial Dirichlet Bootstrap from  Saxena et al.
+(2024), may be necessary.
+7 Disclosure statement
+The authors have the following conflicts of interest to declare.
+8 Data Availability Statement
+All data and code needed to reproduce the analyses in this article are available at https://github.com/Silverman-Lab/MultiAddGPs.
+SUPPLEMENTAL MATERIALS
+Supplement
+includes the following parts: a review of the Collapse-Uncollapsed (CU) sampler and Laplace Approximation in the Collapsed step, a derivation of the posterior distribution for the matrix conjugate linear model, a theoretical result on MLTP for MultiAddGP models, and additional simulation and application details.